### v1

# Smurfs: Leveraging Multiple Proficiency Agents with Context-Efficiency for Tool Planning

Abstract
The emergence of large language models (LLMs) has opened up unprecedented possibilities for automating complex tasks that are often comparable to human performance. Despite their capabilities, LLMs still encounter difficulties in completing tasks that require high levels of accuracy and complexity due to their inherent limitations in handling multifaceted problems single-handedly. This paper introduces "Smurfs", a cutting-edge multi-agent framework designed to revolutionize the application of LLMs. By transforming a conventional LLM into a synergistic multi-agent ensemble, Smurfs enhances task decomposition and execution without necessitating extra training. This is achieved through innovative prompting strategies that allocate distinct roles within the model, thereby facilitating collaboration among specialized agents. The framework gives access to external tools to efficiently solve complex tasks. Our empirical investigation, featuring the mistral-7b-instruct model as a case study, showcases Smurfs’ superior capability in intricate tool utilization scenarios. Notably, Smurfs outmatches the ChatGPT-ReACT in the ToolBench I2 and I3 benchmark with a remarkable 84.4% win rate, surpassing the highest recorded performance of a GPT-4 model at 73.5%. Furthermore, through comprehensive ablation studies, we dissect the contribution of the core components of the multi-agent framework to its overall efficacy. This not only verifies the effectiveness of the framework, but also sets a route for future exploration of multi-agent LLM systems. Our code will be available at https://github.com/FreedomIntelligence/Smurfs/ in the future.
Smurfs: Leveraging Multiple Proficiency Agents with Context-Efficiency for Tool Planning
Junzhi Chen [^1] These authors contributed equally to this work., Juhao Liang [^1] These authors contributed equally to this work., Benyou Wang🖂 Shenzhen Research Institute of Big Data The Chinese University of Hong Kong, Shenzhen wangbenyou@cuhk.edu.cn
footnotetext: 🖂Corresponding author.
Figure 1: Demonstration of the whole process of the Smurfs framework.
1 Introduction
Tool manipulation has traditionally been seen as a distinctive human characteristic, dating back approximately 2.5 million years Oakley and Museum (1972); Ambrose (2001). For large language models (LLMs), access to external tools can equip them with broader capabilities beyond their fixed language modeling knowledge. For example, the search engine API empowers ChatGPT to access real-time information Zhao et al. (2023). However, LLMs still face many challenges when attempting to use tools to solve tasks. These challenges include computational expense and a lack of adaptability to new tools Hao et al. (2024); Guu et al. (2020); Qin et al. (2024).
This paper addresses the critical research problem of enhancing the problem-solving capabilities of LLMs through the adoption of a multi-agent system (MAS) framework Dorri et al. (2018); Van der Hoek and Wooldridge (2008). We posit that a MAS approach can significantly augment the efficacy of LLMs in handling tasks that require a high degree of precision, adaptability, and comprehensive knowledge integration.
ToolBench-all ToolBench-long Increase
ChatGPT-DFSDT 71.5 66.0 $-7.69\%$
GPT4-ReACT 72.0 65.1 $-9.58\%$
GPT4-DFSDT 77.5 69.8 $-9.94\%$
Mistral-Smurfs 79.5 82.1 $+3.27\%$
Table 1: Pass rate of models on the ToolBench benchmark I2 category subset with long settings and all settings.
To this end, we introduce "Smurfs" an innovative MAS framework inspired by the collaborative and versatile nature of its namesake cartoon characters. The Smurfs framework is based on the principle that synergistic collaboration among specialized agents can overcome the limitations faced by individual LLMs. Each agent within the Smurfs framework is designed to perform specific sub-tasks, facilitating a more nuanced and effective approach to complex problem-solving. Our research delves into the architectural design, coordination mechanisms, and the operational dynamics of integrating specialized agents into a cohesive system. Through rigorous experimental evaluation, the Smurfs framework, utilizing the mistral-7b-instruct model Jiang et al. (2023), achieved a remarkable 84.4% win rate against the benchmark set by ChatGPT-ReACT Yao et al. (2022) on the ToolBench I2 and I3 benchmark Qin et al. (2024). This outcome not only sets a new state-of-the-art in the field, but also provides concrete evidence of the effectiveness of the multi-agent approach in enhancing LLM capabilities.
The structure of this paper is as follows: Section 2 presents the motivation for utilizing a multi-agent system. The methodology employed within the framework is detailed in Section 3. Subsequently, Section 4 provides an in-depth evaluation of the experiments conducted on Smurfs. Section 5 then reviews the techniques currently related to our work. Lastly, we summarize and conclude our findings in Section 6.
2 Motivation
2.1 Limited Context Length in a Single Model
LLMs face considerable challenges when tasked with managing extensive contexts. As highlighted by Liu et al. (2024), these limitations become particularly noticeable in tasks requiring assimilation and processing of large inputs, like verbose tool documents and API responses. The situation worsens when LLMs are supplemented with external information, such as document retrieval or online searching Petroni et al. (2020); Ram et al. (2023); Mallen et al. (2022). Although numerous language models capable of handling larger contexts are emerging Dai et al. (2019); Dao et al. (2022), they often face significant performance degradation when the important information is located at some positions Liu et al. (2024); Shi et al. (2023). Moreover, within the MAS framework, the impact of extended contexts on performance remains unclear.
To measure the impact of extended contexts on the performance of LLMs in tool utilization tasks, we conducted a pilot study on the ToolBench benchmark Qin et al. (2024). Additional details can be found in Appendix A.1. We selected the samples with more than 3 steps performed by the ChatGPT-DFSDT method as a subset, called "ToolBench-long", to compare the performance difference with the full set. As demonstrated in Table 1, there is a significant decrease in the pass rate of existing frameworks when faced with tasks involving lengthy questions. This result supports the hypothesis that not only do extended contexts strain the models’ computational efficiency, but they also hinder their ability to accurately interpret and respond to the given instructions. The pilot study highlighted a major issue with current tool utilization frameworks: the excessive context length adversely impacts the models’ planning and execution capabilities when using tools.
Advantages of MAS: The pilot study highlights the need for a new approach beyond the traditional single-agent LLM model. MAS offers a promising solution by distributing tasks among specialized agents. This approach enhances memory efficiency, minimizes distractions, and allows for modular design and optimization of agents. Key benefits include: (1) Memory Efficiency: MAS manages memory better by assigning distinct segments of tasks to different agents, avoiding overload from processing lengthy contexts. (2) Reduced Distractions: Specialized agents focus on specific tasks, reducing interference from irrelevant information and improving overall performance. (3) Modular Design: The modular framework enables individual agent optimization, scalability, and adaptability to diverse tasks.
2.2 Tools Using and Planning in LLMs
Figure 2: A pilot study explores the relationship between dialogue turns and the accuracy of tool selection for LLMs. It demonstrates that long-context instructions undermine the ability of LLMs to select the right tools.
Using external tools in LLMs is a well-established method to enhance model capabilities. However, as the number of available tools grows, managing multiple tools effectively becomes a complex planning challenge Qin et al. (2024). This complexity often results in longer texts for tasks involving multiple tools. In the second pilot study, see details in Appendix A.2, we investigate the relationship between conversation turns and the model’s tool selection accuracy. Figure 2 shows that as the number of dialogue rounds increases, the model’s tool selection accuracy decreases linearly, underscoring the performance impact of longer texts in tool-involved tasks.
While numerous general MAS systems perform well, as shown in studies like Du et al. (2023); Liang et al. (2023), there’s a noticeable absence of multi-agent frameworks specifically designed for tool calling tasks. As highlighted in our two pilot studies, tool usage and planning tasks are particularly sensitive to context length and efficiency. Therefore, there’s a pressing need for an optimized multi-agent framework tailored to tool calling scenarios.
Figure 3: Details of the subtask-solving process of the Smurfs framework. The dotted line represents that the agent can see the memory and the full line stands for operation.
3 Smurfs: A framework with multiple agents
The Smurfs, the beloved cartoon characters, symbolize unity and resourcefulness, and are good at using tools to overcome any challenge they encounter.
3.1 Framework Overview
Figure 1 illustrates the entire workflow for the Smurfs framework. Initially, the Planning Agent identifies the user’s complex request and breaks it down into manageable sub-tasks. Executor Agents are then tasked with collecting this specific information, utilizing access to external tools. Answer Agent compiles the findings into a cohesive response, which is subsequently verified by the Verifier Agent to ensure accuracy and relevance. This process exemplifies the framework’s capability to efficiently handle complex queries by leveraging the specialized roles of multiple agents, thereby ensuring both the precision of task execution and the quality of the output. In the following sections, the functions of each agent will be detailed. More details about the memory of each agent can be seen at Appendix B
Planning Agent
The primary responsibility of the Planning Agent is task decomposition. The strategy known as least-to-most prompting Zhou et al. (2023) is highly effective in dissecting intricate problems into manageable sub-tasks and resolving them sequentially. This approach has demonstrated significant effectiveness and broad applicability. In scenarios involving complex reasoning tasks, we utilize the fundamental principles of the least-to-most strategy to break down intricate tasks into multiple sub-tasks, thereby improving management, efficiency and also the interpretability. An example illustrating how the Planning Agent employs this strategy to decompose a task is provided at Table 2, with the specific prompt available in Appendix C.
Task I’m planning a trip to Turkey and need information about postal codes in Istanbul. Can you provide me with the postal code and district for Istanbul province with plate number 34? Additionally, I would like to know if there are any transit agencies available in Istanbul. Please fetch their names and contact numbers.
Decomposed Sub-tasks 1. Determine the postal code and district for Istanbul province with plate number 34.
2. Find out if there are any transit agencies in Istanbul.
3. Get the names of the transit agencies in Istanbul.
4. Obtain the contact numbers for the transit agencies in Istanbul.
Table 2: Example of task decomposition by Planning Agent.
Executor Agent
The Executor Agent is responsible for choosing and executing the tools to solve the sub-tasks. The agent has access to an external tool library Qin et al. (2024). At each steps, the agent can invoke one tool to tackle the given task. As outlined in Algorithm 1, the agent, using the ReACT format Yao et al. (2022) to choose the tool and arguments, then execute the tool. More detailed information of the Executor Agent can be found in Appendix B.2.
Input: A task $q$, a hint from the verifier agent $h$, problem solving history $H$ and a set of available tools $T$.
Output: Tool response that is useful for solving the task.
thought $\leftarrow$ gen_thought($q$, $h$, $T$);
tool $\leftarrow$ choose_tool($q$, thought, $T$);
args $\leftarrow$ gen_arguments($q$, tool, $H$, tool_doc);
response $\leftarrow$ call_tool(tool, args);
Algorithm 1 Tool Call
Answer Agent
To mitigate the performance degradation caused by lengthy contexts, we introduce the ‘Answer Agent‘ role, designed to extract crucial content for each step and subtask. As demonstrated in the pilot study presented in Section 2.1, retaining all information may not always be beneficial, particularly in cases where the solution path is challenging to discern. Therefore, the primary role of the Answer Agent is to succinctly summarize the generated answers.
Verifier Agent
Similar to current reasoning frameworks Wei et al. (2022); Yao et al. (2022), sequential task reasoning may lead to high computational and tool resource waste within the framework Qin et al. (2024). Therefore, inspired by Qin et al. (2024), we employ a depth-first search with an early stopping strategy to find the solution path. The Verifier Agent serves as an early-stopping and reflection mechanism in the intelligent multi-agent ensemble, allowing for a balance between effectiveness and efficiency. This mechanism not only ensures the accuracy of the generated responses but also optimizes resource use by preventing unnecessary computations. Moreover, if the Verifier Agent thinks the answer at this step isn’t accurate and reasonable enough, it will provide hints for the Executor Agent to get the missing information for the next steps. This dual role of the Verifier Agent enhances the overall performance of the framework, making it more robust and reliable for handling complex tasks.
3.2 Subtask Solving Process
After introducing the function of each agent, this section outlines how the agents collaborate to solve sub-tasks, as shown in Figure 3. Upon receiving a subtask or entering a new step, the tool list is refreshed, allowing the use of all available tools. The Executor Agent then calls the tools according to the task instruction. If a tool call fails, that tool is marked as unusable for this task at this step and the Executor Agent will try using other tools to solve the subtask at this step. This system also introduces a backtracking mechanism similar to DFSDT Qin et al. (2024) to handle the situation where errors frequently occur. Once the executor agent gets a correct response from a tool, the Answer Agent will refine the information, filtering out irrelevant details such as lengthy web pages and generate the answer for the subtask at this step using the local memory. After this, the Verifier Agent checks the answer’s accuracy. If it’s incorrect, the process returns to the Executor Agent with the hint from the Verifier Agent for the next step; Otherwise, the final answer is provided. If the Executor Agent reaches its retry limit without success, Answer Agent will review the entire process from the global memory to produce an answer. More details of the subtask solving process can be seen at  B
4 Experiments
To evaluate the effectiveness and efficiency of the Smurfs framework, we carried out a series of thorough experiments. In addition to the main experiment designed to assess the entire framework, we conducted an ablation study to test the capabilities of each component within the multi-agent framework. This section offers a detailed description of the experimental setup, methodologies employed, and key findings. Our goal is to showcase how the Smurfs framework, through the cooperative work of its multi-agent ensemble, effectively manages complex tasks while optimizing resource use. The experiments were designed with the following research objectives:
  • To validate the capability of the entire framework in managing tool planning tasks;
  • To independently assess the impact of each component on the overall performance, and identify the key factors influencing the multi-agent framework;
4.1 Evaluation
Our experiments are conducted on ToolBench Qin et al. (2024), which encompasses multi-step tool usage tasks across over 16,000 APIs. To evaluate the planning and reasoning capabilities of the LLMs, we focused our experiments on intra-category multi-tool instructions (I2) and intra-collection multi-tool instructions (I3). These instructions involve selecting 2-5 tools from the same category or collection and sampling up to 3 APIs from each tool to formulate the instructions. We employed two metrics for evaluation: (1) Pass Rate measures the percentage of instructions successfully executed within the allocated budget, evaluated by ChatGPT. (2) Win Rate represents the preference selection by a ChatGPT evaluator when presented with two solution paths. All other settings are kept consistent with those of the ToolBench benchmark.
Models Method I2-Inst. I2-Cat. I3-Inst. Average
Pass Win Pass Win Pass Win Pass Win
ToolLLaMA-7B ReACT 30.5 50.8 31.5 41.8 25.0 55.0 29.0 49.2
DFSDT 77.0 68.5 77.0 58.0 66.0 69.0 73.3 65.2
Vicuna-7B ReACT & DFSDT 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Smurfs (ours) 77 73 70.5 64.25 78.0 87.0 75.2 74.8
Mistral-Instruct-7B ReACT & DFSDT 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Smurfs (ours) 77.5 80.0 79.5 79.2 79.0 94.0 78.7 84.4
GPT4 ReACT 67.0 65.8 72.0 60.3 47.0 78.0 62.0 68.0
DFSDT 79.5 73.3 77.5 63.3 71.0 84.0 76.0 73.5
Smurfs (ours) 71.0 77.5 72.0 77.0 64.0 89.5 69.0 81.3
Table 3: ToolBench evaluation, with some results derived from Qin et al. (2024); Yuan et al. (2024). The most effective approach is highlighted in bold, while the second is underlined.
4.2 Baselines
To investigate the varying impacts of the agent framework on models with different capabilities, we categorize our baseline into three groups. The first group consists of models that are fine-tuned based on the tool dataset, represented by ToolLLaMA Qin et al. (2024). The second group encompasses untrained general language models such as Vicuna Chiang et al. (2023), and Mistral-Instruct-7B Jiang et al. (2023). The third category represents the closed-source model, embodied by GPT4. We subsequently contrast our approach with two agent frameworks, ReACT Yao et al. (2022) and DFSDT Qin et al. (2024), both of which are utilized for multi-step reasoning and model invocation. Notably, all methods employ the ground truth toolset for tool selection, thereby eliminating the influence of the tool retriever.
4.3 Main Experiments
Table 3 displays the results of the comprehensive evaluation of our proposed framework on ToolBench. For the untrained LLMs, it is clear that existing agent frameworks do not improve their performance in tool planning tasks; Vicuna, and Mistral-Instruct-7B all failed at the given tasks with the ReACT and DFSDT frameworks. However, Smurfs exhibits exceptional performance: Mistral combined with Smurfs achieves the highest score among the baselines. Through its task decomposition mechanism, Smurfs transforms a long-context task into several simpler tasks, enabling the untrained model to effectively utilize external tools to manage complex tasks. Regarding the closed-source models, specifically GPT4 in these experiments, Smurfs also demonstrates competitive performance on the benchmark compared to other agent frameworks. A high win rate suggests that Smurfs is more adept at finding better solution paths than ChatGPT, a success that can likely be credited to the verifier agent.
4.4 Ablation Study
4.4.1 Importance of each component in MAS
Average
Pass Win
Mistral 0.0 0.0
Mistral with Smurfs 79.3 86.6
   w/o Verifier Agent $74.5_{\downarrow 6.0\%}$ $83.8_{\downarrow 3.2\%}$
   w/o Answer Agent $73.3_{\downarrow 7.6\%}$ $81.7_{\downarrow 5.7\%}$
   w/o Planning Agent $64.0_{\downarrow 19.3\%}$ $82.9_{\downarrow 4.3\%}$
Table 4: Ablation study on ToolBench I2-Cat., I3-Inst. with Mistral-Instruct-7B.
We performed an ablation study to investigate the impact of each agent in our framework. We removed each agent individually, except for the indispensable Executor Agent, and compared the results to the complete framework. Table 4 shows that the Planning Agent is the most crucial component, followed by the Answer Agent and the Verifier Agent. (1) Verifier Agent Removal: Without verification, the framework uses a general depth-first search, leading to increased computational demand and more tool invocations. (2) Answer Agent Removal: Removing this agent means the Executor Agent’s answer won’t be summarized, risking the ’lost-in-the-middle’ problem due to lengthy tool responses. (3) Planning Agent Removal: Without this agent, the global path searching strategy is affected. Models with Smurfs may show reduced performance without preliminary planning, as seen in current frameworks like ReACT and DFSDT.
4.4.2 Effect of number of agents
Models # of I2-Cat. I3-Inst. Average
Agents Pass Win Pass Win Pass Win
Vicuna 1 49.5 53.0 60.0 86.0 54.8 69.5
2 71.0 66.0 75.0 83.0 73.0 74.5
3 70.5 64.3 78.0 87.0 74.3 75.6
Mistral 1 65.5 70.5 56.0 94.0 60.8 82.3
2 76.0 75.5 73.0 92.0 74.5 83.8
3 79.5 79.2 79.0 94.0 79.3 86.6
GPT4 1 60.5 71.8 57.0 89.0 58.8 80.4
2 69.5 77.8 59.0 94.0 64.3 85.9
3 72.0 77.0 64.0 89.5 68.0 83.3
Table 5: The impact of MAS complexity on the performance of different models. More agents don’t always bring more benefits.
Figure 4: The illustration of how GPT4-Smurfs and GPT4-DFSDT solve long context problem. The two sub-questions and their corresponding answers are marked in two colors.
In addition to evaluating the importance of each component, another relevant question arises: Can a fixed agent system effectively accommodate every model? Our hypothesis suggests that larger language models possess comprehensive intelligence, and employing complex agent systems may potentially hinder their performance compared to simpler ones. As shown in Table 5, we conducted an ablation study on Vicuna, Mistral-Instruct-7B, and GPT4, varying the number of agents ***Agents are incrementally added to the system based on their importance.. The findings indicate that as the number of agents in a MAS increases, leading to increased complexity of the MAS, the system’s performance enhancement does not scale linearly. In certain instances, GPT4 with a 2-agent MAS outperforms the one with a 3-agent MAS, suggesting that larger language models do not consistently benefit from more complex Multi-Agent Systems. Smaller language models can benefit more from the agent system, as seen with Vicuna increasing from 69.5 to 75.6.
4.5 Case Study
As shown in Figure 4, even though GPT4-DFSDT and GPT4-Smurfs use the same tool calls to solve the problem, GPT4-DFSDT only answers the first sub-question correctly while GPT4-Smurfs answers both sub-questions accurately. In the process of addressing the second sub-question, it is notable that the tool response only mentions titles of film and television products related to "Star Wars", without addressing OTT platforms. GPT-4-DFSDT erroneously interprets these titles as responses to the question, while GPT-4-Smurfs adeptly identifies this discrepancy and provides a more appropriate response. This case highlights that in situations where tool responses are lengthy and questions are complex, the single agent framework like DFSDT may be susceptible to distractions from irrelevant information, leading to erroneous answers. Conversely, the context-efficient Smurfs framework demonstrates a reduced susceptibility to irrelevant information, thereby generating more accurate answers.
5 Related Work
In this section, we review the literature related to multi-agent collaboration, tool-augmented language models, and complex task planning and reasoning. Each subsection provides an overview of the topic, discusses recent advancements, identifies challenges, and suggests avenues for future research.
5.1 Multi-Agents Collaboration
Multi-agent systems have garnered significant attention due to their applicability in various domains such as robotics, economics, and computer networking. One of the key challenges in multi-agent systems is achieving effective collaboration among autonomous agents. Recent studies have focused on understanding collaborative behaviors and developing coordination strategies to enable agents to work together towards common goals. For example, research in reinforcement learning has explored techniques for emergent coordination, where agents learn to collaborate through interaction with the environment and other agents. Du et al. (2023); Liang et al. (2023) carry out multi-agent interaction in the form of debate, which improves mathematical and strategic reasoning tasks. Li et al. (2023); Wang et al. (2023) uses role-playing to conduct the interaction between multi-agents.
5.2 Tool-Augmented Language Models
Language models augmented with external knowledge sources, often termed as tool-augmented language models, have demonstrated promising outcomes in a variety of natural language processing tasks. These models utilize external knowledge bases, ontologies, or pre-trained models to enhance their contextual understanding and boost performance in tasks such as text generation, summarization, and question answering. Recent advancements in this field include techniques for integrating structured knowledge into language models, such as graph-based representations or semantic parsing. For instance, Qin et al. (2024) introduces ToolBench, an instruction-tuning dataset for tool usage, along with a fine-tuned tool-oriented model, ToolLLaMA. Another significant contribution is Gorilla Patil et al. (2023), which excels at writing API calls and also introduces a benchmark for evaluating LLMs with tools.
5.3 Complex task planning and reasoning
Problem decomposition (Zhou et al. (2023); Drozdov et al. (2022); Khot et al. (2022); Press et al. (2022)) is a popular paradigm used in the LLM and leads to good performance in challenging reasoning tasks. This divide-and-conquer strategy provides sufficient explanation of how the model works. The current trend is that when model size reaches some limit, people start looking for a way to make the most of the limited model. Chain-of-thought prompting Wei et al. (2022) is one of the strategy that lead models to divide the tasks into a chain, and solve the task step by steps. Chen et al. (2022) propose a program of thoughts structure to enhance the reasoning ability. Yao et al. (2024) simulate thought as tree structure to solve task. Besta et al. (2024) using graph to simulation.
6 Conclusion
In this study, we present a novel MAS framework, Smurfs, tailored to enhance the planning and reasoning capabilities of LLMs in handling complex tasks that involve lengthy contexts and tools. We conducted experiments on the multi-step tool usage benchmark, ToolBench, and the results demonstrated the overall effectiveness of the Smurfs framework compared to the baseline models.
Ablation studies are carried out to investigate and compare the significance of different components within the MAS framework. The findings revealed that preliminary planning was the most crucial element. Content summarizing also played a key role in mitigating the ’lost-in-the-middle’ issue often encountered in long-context multi-step reasoning scenarios. While verification was not as influential on effectiveness, it proved valuable in enhancing computational efficiency by identifying the optimal solution path for complex task resolution. By dissecting and comparing the different aspects of the MAS framework, we aim to offer insights that could inspire advancements in the applicability and accuracy of LLMs.
In conclusion, this research contributes to the expanding field of study focused on enhancing LLM capabilities, particularly for multi-step tool usage tasks. It emphasizes the importance of task decomposition, preliminary planning, and efficient verification for improving task execution performance. We are confident that the knowledge gained from this study will lay the groundwork for the development of more sophisticated and efficient LLM frameworks in the future.
7 Limitations
Generalization Ability:
While our empirical investigation has demonstrated promising results for the Mistral-7b-instruct model, additional benchmark evaluations may be needed to validate the generalization ability of the proposed Smurfs framework across various LLM architectures and tasks.
Model Size Constraints:
Due to device limitations and computational constraints, our experiments primarily focused on the 7B models. Further evaluations with larger and smaller LLMs are required to assess the impact of the Smurfs framework on models of different sizes.
Computational Efficiency:
Although preliminary findings suggest that the Smurfs framework can enhance computational efficiency through efficient verification, a more detailed analysis is needed to quantify the computational overhead introduced by the multi-agent architecture.
Acknowledging these limitations, future research should aim to address these gaps to provide a more comprehensive understanding of the Smurfs framework’s capabilities and potential areas for improvement.
Appendix A Pilot Study Settings
Figure 5: Demonstration of the memory of the Smurfs framework.
A.1 Pilot Study 1
We conducted the first pilot study using the ToolBench benchmark Qin et al. (2024), which comprises over 16,000 APIs focused on tool usage and planning. We divided the benchmark into two subsets: "ToolBench-long," containing samples requiring more than three steps for ChatGPT-DFSDT, and "ToolBench-all," which includes all samples from ToolBench. On average, we consider the "ToolBench-long" subset to be more complex than the "ToolBench-all" set. We maintained all other settings consistent with the original source settings.
A.2 Pilot Study 2
The second pilot study is conducted using the BMTools Qin et al. (2023) test datasets. We augmented each sample by adding random, irrelevant test samples as prefix prompts to increase their input length. We used tool selection accuracy as the metric to evaluate the LLMs’ model selection capabilities. The experiment uses Mistral-7B-Instruct-v0.1 Jiang et al. (2023) and Qwen1.5-7B-Chat Bai et al. (2023) for evaluation.
Appendix B Details of the Smurfs
Figure 6: Details of the executor agent working process
B.1 Memory management of Smurfs
When each agent performs a task, the prompt information they receive varies. We refer to this as the memory management system of Smurfs. As illustrated in Figure 5, information is primarily divided into four categories. The first category is local memory, consisting of the thought and answer from the previous steps. When backtracking happens, the thought and answer of the backtracking step will be popped out from the local memory. The second category is global memory, which stores all the previous steps history including those that have been backtracked. The third category is the tool list, similar to other tool-using frameworks, storing brief information about all accessible tools. Lastly, the tool document provides detailed usage information like the parameters infomation for each tool listed.
As illustrated in Figure 3, each agent has its own memory context. The Executor Agent’s memory will be discussed in the next section. The Answer Agent takes in the subtask, the tool response and the local memory to generate the answers to the subtask using all the local memory so far. Then the thought and answer of this step will be added to the local memory and the global memory. The Verifier Agent takes in the subtask, the answer from the Answer Agent and return the status of the subtask. If the status is solved, then the system will return the subtask’s answer and process to the next subtask; If the status is pending, the Verifier Agent will give hint to the Executor Agent to process to the next step. When the max steps have been reached for a subtask, the answer agent will generate the answer for the subtask using the global memory and process to the next subtask.
B.2 Executor Agent Workflow
Next, let’s illustrate how memory is accessed during the Executor Agent’s working process, as shown in Figure 6. When a subtask arises, the Executor Agent first receives the hint (task instruction) from the Verifier Agent, the subtask, the available tool list to generate the thought(a strategy for task execution) at this step. It then takes in the subtask, the thought and the tool list to generate an action with a specific action name. Subsequently, it takes in the subtask, the thought, the tool document of the chosen action and the local memory to generate the input of the chosen action.
B.3 Backtracking Mechanism
Smurfs uses a backtracking mechanism similar to DFSDT  Qin et al. (2024). As illustrated in Figure 3, when tool responses error, the system will delete the tool from the tool list and retry this step. However, when all tools have reported error or the model thinks the available tools can’t solve the subtask at this step, the system will pop out the latest local memory and the latest step, delete the tool used by the popped out step from the tool list of the current step and retry the the other solution path.
Appendix C Prompts for multi-agent implementation
Prompts used by each agent and their example outputs are shown in Figure  7 to  13.
Planning Agent Prompt: You need to decompose a complex user’s question into some simple sub-tasks and let the model execute it step by step. Please note that: 1. You should only decompose this complex user’s question into some simple sub-tasks which can be executed easily by using a single tool. 2. Each simple subtask should be expressed into natural language. 3. Each subtask should contain the necessary information from the original question and should be complete, explicit and self-consistent. 4. You must ONLY output in a parsible JSON format. An example output looks like: ”’ {"Tasks": ["Task 1", "Task 2", …]} ”’ This is the user’s question: I’m planning a trip to Turkey and need information about postal codes in Istanbul. Can you provide me with the postal code and district for Istanbul province with plate number 34? Additionally, I would like to know if there are any transit agencies available in Istanbul. Please fetch their names and contact numbers. Output: "Tasks": ["Find the postal codes and districts for plate number 34 in Istanbul.", "Search for transit agencies and their contact numbers in Istanbul."] This is the user’s question: I recently moved to a new address and I need to update my information. Can you retrieve my address details using the postal code 75094080? Additionally, I would like to know the companies that offer shipping services. Output: {"Tasks": ["retrieve the address details using the postal code 75094080", "search for companies that offer shipping services to my address"]} This is the user’s question: {question} Output: Example Output: {"Tasks": ["Determine the postal code and district for Istanbul province with plate number 34.", "Find out if there are any transit agencies in Istanbul.", "Get the names of the transit agencies in Istanbul.", "Obtain the contact numbers for the transit agencies in Istanbul."] }
Figure 7: An example prompt for task decomposition in the framework.
Tool Check Prompt: As a powerful language model, you’re equipped to answer user’s question with accumulated knowledge. However, in some cases, you need to use external APIs to answer accurately. Thus, you need to check whether the user’s question requires you to call an external API to solve it. Here are some tips to help you check: 1. If the user’s question requires real-time information, since your knowledge base isn’t updated in real-time, any such question will demand an API call. 2. If you need to obtain information (e.g., ID, name, phone number, geographical location, rank, etc.), you need to call the database APIs if you are not sure. 3. If the question demand a database search or internet research to generate an answer, this is another situation where an API call is necessary. If need, please output ’YES’; If not, please output ’NO’ You need to give reasons first and then decide whether to keep it or not. You must only output in a parsible JSON format. Two example outputs look like: Example 1: {"Reason": "The reason why you think you do not need to call an external API to solve the user’s question", "Choice": "No"} Example 2: {"Reason": "The reason why you think you need to call an external API to solve the user’s question", "Choice": "Yes"} This is the user’s question: {question} Output: Example Output: {"Reason": "To determine the postal code and district for a specific location based on a plate number, we would typically need to access a combination of databases, including vehicle registration databases and postal code databases. Since we do not have direct access to these databases, we will need to call external APIs to retrieve this information. Therefore, the user’s question requires an API call.", "Choice": "Yes"}
Figure 8: An example prompt for tool check in the framework.
Executor Agent: Thought Prompt: You need to analyse the previous execution history and generate your internal reasoning and thoughts on the task, and how you plan to solve it based on the current attempts. Do not output thought that is too long. Output in 2-3 sentences is OK. This is the user’s task: {task} This is the Tool List: {functions} This is the previous execution history: {messages} This is the hint comes from the evaluator: {hint} Output: Example Output: Looking at the task, the user wants to find the postal codes and districts for the number plate 34 in Istanbul. I have a tool that provides Turkish plates, and since Istanbul is in Turkey, this tool might provide the needed information. I will proceed to use the ’Logistics:Turkey Postal Codes:il’ tool to accomplish the task.
Figure 9: An example prompt for tool check in the framework.
Executor Agent: Action Prompt: This is the user’s question: {question} These are the tools you can select to solve the question: Tool List: {Tool_list} Please note that: 1. You should only chooce one tool from the Tool List to solve this question. 2. You must ONLY output the ID of the tool and your reason for choosing it in a parsible JSON format. An example output looks like: ”’ Example: {"ID": ID of the tool, "Reason": The reason for choosing the tool} ”’ Output: Example Output: {’ID’: ’1’, ’Reason’: ’This tool provides information about Turkish plates and postal codes, which can be applied to the given task of finding the postal codes and districts for the number plate 34 in Istanbul.’}
Figure 10: An example prompt for action generation in the framework.
Executor Agent: Action Input Prompt: Given a user’s question and a API tool documentation, you need to output parameters according to the API tool documentation to successfully call the API to solve the user’s question. Please note that: 1. The Example in the API tool documentation can help you better understand the use of the API. 2. Ensure the parameters you output are correct. The output must contain the required parameters, and can contain the optional parameters based on the question. If no paremters in the required parameters and optional parameters, just leave it as {} 3. If the user’s question mentions other APIs, you should ONLY consider the API tool documentation I give and do not consider other APIs. 4. The question may have dependencies on answers of other questions, so we will provide logs of previous questions and answers for your reference. 5. You must ONLY output in a parsible JSON Format. The example output looks like: ”’ Example: {"keyword": "Artificial Intelligence", "language": "English"} ”’ There are logs of previous questions and answers: previous_log This is the current user’s question: question This is API tool documentation: api_dic Output: Example Output: {"il": 34}
Figure 11: An example prompt for action input generation in the framework.
Answer Agent Prompt: You should answer the question based on the response output by the API tool. Please note that: 1. Answer the question in natural language based on the API response reasonably and effectively. 2. The user cannot directly get API response, so you need to make full use of the response and give the information in the response that can satisfy the user’s question in as much detail as possible. 3. Do not output answer that is too long. Output in 3-6 sentences is OK. This is the user’s question: {question} This is the API response: {call_result} Output: Example Output: The postal codes and districts for the plate number 34 in Istanbul include: Adalar district with postal code 34975 for Burgazada area, 34970 for Büyükada area, 34973 for Heybeliada area and 34977 for Kınalıada area. There is also the Arnavutköy district with postal code 34275 for areas like Anadolu, Arnavutköy Merkez, İmrahor, İslambey, Mustafa Kemal Paşa, Nenehatun, and Yavuz Selim. Another area in Arnavutköy district is Baklalı with postal code 34277.
Figure 12: An example prompt for Answer Agent in the framework.
Verifier Agent Prompt: An agent is trying to solve the query proposed by the user. You need to evaluate whether the given query has been completed reasonably and accurately. If so, summarize the solution to the user. If not, summarize the current progress, and propose what is missing. You response contains following elements: Speak: (your words to the agent if the task is pending, or a complete answer based on the full execution log to the user if the task is finished) Status: (0 or 1. 0 for pending and 1 for finished) You must only output in a parsible JSON format. Two example outputs look like: Example 1: {"Speak": "answer based on the full execution log to the user", "Status": "1"} Example 2: {"Speak": "your words to the group if the task is pending", "Status": "0"} This is the answer from the previous execution result: {answer} This is the original question: {question} Output: Example Output: {"Speak": "The postal codes and districts for plate number 34 in Istanbul include: Adalar district with postal code 34975 for Burgazada area, 34970 for Büyükada area, 34973 for Heybeliada area and 34977 for Kınalıada area. In addition, Arnavutköy district has postal code 34275 for areas such as Anadolu, Arnavutköy Merkez, İmrahor, İslambey, Mustafa Kemal Paşa, Nenehatun, and Yavuz Selim. Another part of Arnavutköy district, Baklalı, has the postal code 34277.", "Status": "1"}
Figure 13: An example prompt for Verifier Agent in the framework.

### v3

@@
 Despite their capabilities, LLMs still encounter difficulties in completing tasks that require high levels of accuracy and complexity due to their inherent limitations in handling multifaceted problems single-handedly.
-This paper introduces "Smurfs", a cutting-edge multi-agent framework designed to revolutionize the application of LLMs.
-By transforming a conventional LLM into a synergistic multi-agent ensemble, Smurfs enhances task decomposition and execution without necessitating extra training.
-This is achieved through innovative prompting strategies that allocate distinct roles within the model, thereby facilitating collaboration among specialized agents.
-The framework gives access to external tools to efficiently solve complex tasks.
-Our empirical investigation, featuring the mistral-7b-instruct model as a case study, showcases Smurfs’ superior capability in intricate tool utilization scenarios.
-Notably, Smurfs outmatches the ChatGPT-ReACT in the ToolBench I2 and I3 benchmark with a remarkable 84.4% win rate, surpassing the highest recorded performance of a GPT-4 model at 73.5%.
+This paper introduces ‘Smurfs’, a cutting-edge multi-agent framework designed to revolutionize the application of LLMs.
+By seamlessly transforming a conventional LLM into a synergistic multi-agent ensemble, Smurfs can enhance the model’s ability to solve complex tasks at no additional cost.
+This is achieved through innovative prompting strategies that allocate distinct roles within the model, thereby facilitating collaboration among specialized agents and forming an intelligent multi-agent system.
+Our empirical investigation on both open-ended task of StableToolBench and closed-ended task on HotpotQA showcases Smurfs’ superior capability in intricate tool utilization scenarios.
+Notably, Smurfs outmatches all the baseline methods in both experiments, setting new state-of-the-art performance.
 Furthermore, through comprehensive ablation studies, we dissect the contribution of the core components of the multi-agent framework to its overall efficacy.
 This not only verifies the effectiveness of the framework, but also sets a route for future exploration of multi-agent LLM systems.
-Our code will be available at https://github.com/FreedomIntelligence/Smurfs/ in the future.
+Our code is available at https://github.com/FreedomIntelligence/Smurfs/.
 Smurfs: Leveraging Multiple Proficiency Agents with Context-Efficiency for Tool Planning
@@
 (2023).
-However, LLMs still face many challenges when attempting to use tools to solve tasks.
-These challenges include computational expense and a lack of adaptability to new tools Hao et al.
+However, LLMs still encounter several challenges when using multiple tools to solve tasks.
+These challenges include effective solution planning and adaptability to new tools.
+Hao et al.
 (2024); Guu et al.
@@
 (2024).
-This paper addresses the critical research problem of enhancing the problem-solving capabilities of LLMs through the adoption of a multi-agent system (MAS) framework Dorri et al.
+This paper addresses the critical research problem of enhancing the problem-solving capabilities of LLMs through the adoption of a plug-and-play multi-agent system (MAS) framework Dorri et al.
 (2018); Van der Hoek and Wooldridge (2008).
 We posit that a MAS approach can significantly augment the efficacy of LLMs in handling tasks that require a high degree of precision, adaptability, and comprehensive knowledge integration.
-ToolBench-all ToolBench-long Increase
-ChatGPT-DFSDT 71.5 66.0 $-7.69\%$
-GPT4-ReACT 72.0 65.1 $-9.58\%$
-GPT4-DFSDT 77.5 69.8 $-9.94\%$
-Mistral-Smurfs 79.5 82.1 $+3.27\%$
-Table 1: Pass rate of models on the ToolBench benchmark I2 category subset with long settings and all settings.
-To this end, we introduce "Smurfs" an innovative MAS framework inspired by the collaborative and versatile nature of its namesake cartoon characters.
-The Smurfs framework is based on the principle that synergistic collaboration among specialized agents can overcome the limitations faced by individual LLMs.
+Pass Rate $\uparrow$ (%) Win Rate $\uparrow$ (%) # of Tokens per request $\downarrow$ # of Tokens per query $\downarrow$
+ReACT $44.4_{\pm 1.1}$ base 1,424 6,479
+DFSDT $55.4_{\pm 2.0}$ 60.4 1,743 20,714
+Smurfs (ours) $57.4_{\pm 1.1}$ 62.4 459 8,096
+Table 1: Comparison of token cost and performance between tool planning methods over StableToolBench.
+Existing methods, ReACT and DFSDT, have limitations due to high token costs or poor performance.
+The results are averaged over the subtasks within StableToolBench.
+To this end, we introduce ‘Smurfs’ an innovative MAS framework inspired by the collaborative and versatile nature of its namesake cartoon characters.
+The proposed framework is based on the principle: synergistic collaboration among specialized agents can overcome the limitations faced by individual LLMs.
 Each agent within the Smurfs framework is designed to perform specific sub-tasks, facilitating a more nuanced and effective approach to complex problem-solving.
 Our research delves into the architectural design, coordination mechanisms, and the operational dynamics of integrating specialized agents into a cohesive system.
-Through rigorous experimental evaluation, the Smurfs framework, utilizing the mistral-7b-instruct model Jiang et al.
-(2023), achieved a remarkable 84.4% win rate against the benchmark set by ChatGPT-ReACT Yao et al.
-(2022) on the ToolBench I2 and I3 benchmark Qin et al.
+The effectiveness of Smurfs is validated through both open-ended and closed-ended tool planning benchmark experiments Guo et al.
+(2024); Yang et al.
+(2018), where the proposed MAS system consistently outperform baseline methods on both benchmarks.
+An ablation study followed by a case study further investigates the underlying reasons for this effectiveness.
+These results not only establish a new state-of-the-art in the field but also offer concrete evidence of the multi-agent approach’s efficacy in enhancing LLM capabilities.
+The contributions of this paper can be summarized as follows:
+  1. 1.
+    We introduce a novel plug-and-play MAS framework to enhance the tool planning capabilities of LLMs.
+Experiments demonstrate the effectiveness of this approach, which is also more cost-efficient compared to existing tool planning methods.
+  2. 2.
+    Ablation studies further reveal the underlying reasons for the effectiveness of the MAS framework, providing valuable insights for future research.
+2 Motivation
+2.1 Multi-Tool Planning
+To augment LLMs to do multi-tool planning for solving complex problems, previous work has seen numerous attempts.
+Chain-of-Thought  Wei et al.
+(2023) was the first to propose the method of thought and answer chain reasoning.
+ReACT Yao et al.
+(2022) further introduced the thought-action-observation format for tool chain reasoning, leading to the development of various multi-tool planning methods Chen et al.
+(2023a); Xu et al.
+(2023); Shinn et al.
 (2023).
-This outcome not only sets a new state-of-the-art in the field, but also provides concrete evidence of the effectiveness of the multi-agent approach in enhancing LLM capabilities.
-The structure of this paper is as follows: Section 2 presents the motivation for utilizing a multi-agent system.
-The methodology employed within the framework is detailed in Section 3.
-Subsequently, Section 4 provides an in-depth evaluation of the experiments conducted on Smurfs.
-Section 5 then reviews the techniques currently related to our work.
-Lastly, we summarize and conclude our findings in Section 6.
-2 Motivation
-2.1 Limited Context Length in a Single Model
-LLMs face considerable challenges when tasked with managing extensive contexts.
+The latest work, DFSDT Qin et al.
+(2024), was proposed to address the inherent limitations of CoT and ReACT: error propagation and limited exploration.
+Deep First Search Decision Tree, denoted as DFSDT, is powerful in addressing multi-tool planning problems.
+Its core concept involves employing a depth-first search (DFS) approach for multi-tool planning (for more details, see Appendix A).
+When a tool fails or is deemed inadequate for solving the current problem, DFSDT backtracks to the previous solution state and attempts to resolve the issue using a different tool.
+However, several limitations were identified with the mechanism of DFSDT: (1) instability of the rollback mechanism, (2) redundant context, and (3) premature termination.
+The following sections will introduce these limitations in detail.
+2.1.1 Instability of the Rollback Mechanism
+The rollback mechanism in DFSDT is determined by the model.
+The number of steps to roll back and the selection of new tools after rollback are guided using prompt containing the errors encountered in the previous failed trajectory.
+When the model is sufficiently robust, this rollback mechanism serves as a highly flexible and efficient planning strategy.
+However, when the model’s capability is insufficient, it will fail to execute the correct rollback mechanism, i.e.
+retry the same error tools or roll back too far.
+2.1.2 Redundant Context
+In the process of planning with DFSDT, each tool plan is generated using the entire conversation history (including all the thoughts, actions, action inputs and tool responses) as context.
+However, in reality, each step of tool planning only requires a very small portion of the relevant history for effective planning.
+The context redundancy not only increases computational overhead but also reduces the accuracy of model inference due to the inclusion of irrelevant historical data.
 As highlighted by Liu et al.
-(2024), these limitations become particularly noticeable in tasks requiring assimilation and processing of large inputs, like verbose tool documents and API responses.
+(2024), redundant context become particularly noticeable in tasks requiring assimilation and processing of large inputs, like verbose tool documents and API responses.
 The situation worsens when LLMs are supplemented with external information, such as document retrieval or online searching Petroni et al.
@@
 (2024); Shi et al.
-(2023).
-Moreover, within the MAS framework, the impact of extended contexts on performance remains unclear.
-To measure the impact of extended contexts on the performance of LLMs in tool utilization tasks, we conducted a pilot study on the ToolBench benchmark Qin et al.
-(2024).
-Additional details can be found in Appendix A.1.
-We selected the samples with more than 3 steps performed by the ChatGPT-DFSDT method as a subset, called "ToolBench-long", to compare the performance difference with the full set.
-As demonstrated in Table 1, there is a significant decrease in the pass rate of existing frameworks when faced with tasks involving lengthy questions.
-This result supports the hypothesis that not only do extended contexts strain the models’ computational efficiency, but they also hinder their ability to accurately interpret and respond to the given instructions.
-The pilot study highlighted a major issue with current tool utilization frameworks: the excessive context length adversely impacts the models’ planning and execution capabilities when using tools.
-Advantages of MAS: The pilot study highlights the need for a new approach beyond the traditional single-agent LLM model.
-MAS offers a promising solution by distributing tasks among specialized agents.
-This approach enhances memory efficiency, minimizes distractions, and allows for modular design and optimization of agents.
-Key benefits include: (1) Memory Efficiency: MAS manages memory better by assigning distinct segments of tasks to different agents, avoiding overload from processing lengthy contexts.
-(2) Reduced Distractions: Specialized agents focus on specific tasks, reducing interference from irrelevant information and improving overall performance.
-(3) Modular Design: The modular framework enables individual agent optimization, scalability, and adaptability to diverse tasks.
-2.2 Tools Using and Planning in LLMs
-Figure 2: A pilot study explores the relationship between dialogue turns and the accuracy of tool selection for LLMs.
-It demonstrates that long-context instructions undermine the ability of LLMs to select the right tools.
-Using external tools in LLMs is a well-established method to enhance model capabilities.
-However, as the number of available tools grows, managing multiple tools effectively becomes a complex planning challenge Qin et al.
-(2024).
-This complexity often results in longer texts for tasks involving multiple tools.
-In the second pilot study, see details in Appendix A.2, we investigate the relationship between conversation turns and the model’s tool selection accuracy.
-Figure 2 shows that as the number of dialogue rounds increases, the model’s tool selection accuracy decreases linearly, underscoring the performance impact of longer texts in tool-involved tasks.
-While numerous general MAS systems perform well, as shown in studies like Du et al.
-(2023); Liang et al.
-(2023), there’s a noticeable absence of multi-agent frameworks specifically designed for tool calling tasks.
-As highlighted in our two pilot studies, tool usage and planning tasks are particularly sensitive to context length and efficiency.
-Therefore, there’s a pressing need for an optimized multi-agent framework tailored to tool calling scenarios.
+(2023), which is known as the ‘lost-in-the-middle’ problem.
+2.1.3 Premature Termination
+The termination mechanism set by DFSDT involves adding a termination tool to the model’s selectable toolkit.
+When the model selects this termination tool, DFSDT stops and provides an answer.
+However, in practical applications, this mechanism often prematurely terminates when dealing with complex problems requiring multi-step reasoning.
+We hypothesize that this issue arises due to the redundant interference of other tool information and history information, which disrupts the model’s ability to judge whether the original problem should be terminated.
+Instead, the model focuses on whether the current sub-problem requires termination, leading the mechanism to terminate after resolving the sub-problem.
+2.2 Multi Agent System
+Method Multi-Agent Training Generality Reflection Planning
+  ReAct Yao et al. (2022) Iterative
+Reflexion Shinn et al. (2023) Iterative
+Chameleon Lu et al. (2023) Global
+HuggingGPT Shen et al. (2023) Global
+BOLAA Liu et al. (2023) Iterative
+AgentVerse Chen et al. (2023b) Iterative
+FireAct Chen et al. (2023a) Iterative
+DFSDT Qin et al. (2024) Iterative
+RestGPT Song et al. (2023) Iterative
+Lumos Yin et al. (2024) Iterative or Global
+AutoAct Qiao et al. (2024) Iterative
+Smurfs (Ours) Iterative and Global
+Table 2: Comparison of related works.
+To address the limitations inherent in DFSDT and to further enhance LLM’s multi-tool planning capabilities, multi-agent system (MAS) has emerged as a natural solution.
+Inspired by human social division of labor and cooperation, MAS aim to enable AI agents to accomplish more complex tasks more effectively and efficiently through the division of labor and collaboration.
+Previous works Song et al.
+(2023); Liu et al.
+(2023); Chen et al.
+(2023b); Yin et al.
+(2024); Qiao et al.
+(2024) has leveraged MAS to achieve this goal.
+Table 2 shows the difference between them.
+Based on those works, we further design the MAS named Smurfs to address issues with DFSDT.
+By dividing tasks among different agents, each agent can focus on a specific part of the DFSDT task, accessing only the necessary history as context during task execution, which effectively addresses the issue of redundant context.
+The redesign of the rollback mechanism to incorporate memory and tool list rollback mechanisms addresses the instability of the rollback mechanism.
+Drawing on the concept of least-to-most prompting Zhou et al.
+(2023), the original problem is first decomposed into sub-problems for macro-level planning.
+Subsequently, DFSDT is used to solve each sub-problem at the micro-level, with macro-level planning guiding the micro-level planning, thereby resolving the issue of premature termination.
 Figure 2: Details of the subtask-solving process of the Smurfs framework. The dotted line represents that the agent can see the memory and the full line stands for operation.
@@
 Initially, the Planning Agent identifies the user’s complex request and breaks it down into manageable sub-tasks.
-Executor Agents are then tasked with collecting this specific information, utilizing access to external tools.
+Executor Agents are then tasked with collecting task specific information, utilizing access to external tools.
 Answer Agent compiles the findings into a cohesive response, which is subsequently verified by the Verifier Agent to ensure accuracy and relevance.
+Each agent focus on its own task and only use the relevant part of the conversation history to reduce the Redundant Context.
 This process exemplifies the framework’s capability to efficiently handle complex queries by leveraging the specialized roles of multiple agents, thereby ensuring both the precision of task execution and the quality of the output.
-In the following sections, the functions of each agent will be detailed.
-More details about the memory of each agent can be seen at Appendix B
+In the following sections, the system mechanism and functions of each agent will be detailed.
+More details of memory system can be seen at Appendix B.
+3.2 Agent Components
+Tools
+The tool documents about the tools that Smurfs can utilize in the completion of a complex task are denoted as $D=\{n_{i},d_{i},p_{i}\}^{|d|}_{i=1}$, where n represents the tool name, d represents tool usage description, p represents parameter description and $|d|$ represents the amount of the available tools.
+The available tool list is denoted as $\tau=\{n_{i},d_{i}\}^{|\tau|}_{i=1}$.
+$\tau_{t}$ denotes the tool list Smurfs can utilize at time t.
+Memory
+The memory of the agent system at time t is the history of the task-solving process before t, denoted as $M=(m_{1},m_{2},...,m_{t-1})$ and $m_{i}=(\gamma_{i},a_{i})$, where $m_{i}$ represents memory element at time i and $\gamma_{i}$, $a_{i}$ represents thought and answer generated by the system at time i.
+There are two types of memory in Smurfs: local memory and global memory.
+the local memory is used to record the ongoing solution trajectory and to generate the next action in the current trajectory.
+The global memory, meanwhile, records all trajectories and is used to generate the sub-problem’s answer by combining all trajectory records when the maximum number of retries is exceeded.
+This local-global combined memory system ensures that the planning of the current solution trajectory is not influenced by the context of erroneous trajectories.
+It also generates an answer that combines all trajectories when the verifier agent cannot determine task completion within the maximum number of planned steps.
+This memory system ensures context efficiency during the task-solving process.
+3.3 Macro Planning
 Planning Agent
-The primary responsibility of the Planning Agent is task decomposition.
-The strategy known as least-to-most prompting Zhou et al.
-(2023) is highly effective in dissecting intricate problems into manageable sub-tasks and resolving them sequentially.
-This approach has demonstrated significant effectiveness and broad applicability.
-In scenarios involving complex reasoning tasks, we utilize the fundamental principles of the least-to-most strategy to break down intricate tasks into multiple sub-tasks, thereby improving management, efficiency and also the interpretability.
-An example illustrating how the Planning Agent employs this strategy to decompose a task is provided at Table 2, with the specific prompt available in Appendix C.
-Task I’m planning a trip to Turkey and need information about postal codes in Istanbul.
-Can you provide me with the postal code and district for Istanbul province with plate number 34?
-Additionally, I would like to know if there are any transit agencies available in Istanbul.
-Please fetch their names and contact numbers.
-Decomposed Sub-tasks 1. Determine the postal code and district for Istanbul province with plate number 34.
-2. Find out if there are any transit agencies in Istanbul.
-3. Get the names of the transit agencies in Istanbul.
-4. Obtain the contact numbers for the transit agencies in Istanbul.
-Table 2: Example of task decomposition by Planning Agent.
+The primary responsibility of the Planning Agent is doing macro-level planning through task decomposition to prevent premature termination. The inference process of the Planning Agent is:
+$Plan\>P:(p1,p2,...)=PA(q)$ (1)
+Where $p_{i}$ represents sub-problem of the original query q, PA represents the Planning Agent.
+After the task decomposition, the agent system will use Executor Agent, Answer Agent an Verifier Agent to solve each sub-problem using DFSDT collaboratively in a sequential order.
+To utilize the answer of the previous sub-problem when solving subsequent sub-problem, the strategy known as least-to-most prompting Zhou et al.
+(2023) is used.
+3.4 Subtask Solving Process
+After introducing the function of plan agent, this section outlines how the agents collaborate to solve sub-tasks, as shown in Figure 2.
+Stable Rollback
+To address the instability of the rollback mechanism in DFSDT, we propose a rollback mechanism based on rules.
+Whenever an error occurs while using a tool $\tau_{t,i}$ at time t, the tool list at t $\tau_{t}$ will pop $\tau_{t,i}$ out and reperform tool selection and tool planning (ensuring that the faulty tool is not selected again).
+If, at time t, the tool list becomes empty, it signifies that after the system choosing tool $\tau_{t-1,j}$ at time t-1, no subsequent trajectory can solve the problem.
+In this case, the agent system will roll back to time t-1, meaning that the local memory M will pop out the memory element $m_{t-1}$ at time t-1, and the tool list at time t-1 $\tau_{t-1}$ will pop out tool $\tau_{t-1,j}$.
+The agent system will then set the time t=t-1 and continue planning.
+This rule-based rollback mechanism, compared to the original model-based rollback mechanism of DFSDT, is less flexible and might reduce rollback efficiency.
+However, it is more stable, ensuring the correctness of deep first search and enabling models with weaker capabilities to utilize DFSDT on tool planning.
+Backbone Method StableToolBench
+I1-Inst. I1-Cat. I1-Tool. I2-Cat. I2-Inst. I3-Inst. Average
+Pass Win Pass Win Pass Win Pass Win Pass Win Pass Win Pass Win
+GPT-3.5 Turbo ReACT $41.6_{\pm 1.2}$ / $48.4_{\pm 0.5}$ / $52.5_{\pm 0.5}$ / $52.2_{\pm 1.0}$ / $31.6_{\pm 1.2}$ / $39.9_{\pm 2.0}$ / $44.4_{\pm 1.1}$ ///
+GPT-3.5 Turbo DFSDT $54.1_{\pm 1.0}$ $64.4$ $60.1_{\pm 0.0}$ $61.4$ $59.9_{\pm 1.7}$ $53.8$ $60.9_{\pm 0.9}$ $62.9$ $52.8_{\pm 3.7}$ $66.0$ $44.3_{\pm 4.8}$ $54.1$ $55.4_{\pm 2.0}$ $60.4$
+GPT-3.5 Turbo Smurfs $\underline{60.3_{\pm 1.5}}$ $65.0$ $67.0_{\pm 1.0}$ $\underline{69.9}$ $60.3_{\pm 1.3}$ $54.4$ $54.3_{\pm 0.4}$ $\underline{63.7}$ $42.6_{\pm 1.6}$ $64.2$ $60.1_{\pm 1.0}$ $57.4$ $57.4_{\pm 1.1}$ $62.4$
+Mistral-7B ReACT 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
+Mistral-7B DFSDT 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
+Mistral-7B Smurfs $\mathbf{76.3_{\pm 0.8}}$ 63.8 $\mathbf{86.7_{\pm 1.2}}$ 62.7 $\mathbf{81.0_{\pm 1.9}}$ 58.2 $\mathbf{70.4_{\pm 2.7}}$ 54.0 $\mathbf{63.8_{\pm 2.4}}$ $\underline{67.0}$ $\mathbf{85.2_{\pm 0.7}}$ $57.4$ $\mathbf{77.2_{\pm 1.6}}$ $60.5$
+GPT-4 Turbo ReACT $41.1_{\pm 1.5}$ $60.1$ $53.2_{\pm 1.3}$ $62.1$ $42.2_{\pm 1.1}$ $48.1$ $50.0_{\pm 0.7}$ $57.3$ $38.7_{\pm 0.8}$ $65.1$ $37.7_{\pm 1.3}$ $47.5$ $43.8_{\pm 1.1}$ $56.7$
+GPT-4 Turbo DFSDT $52.7_{\pm 1.4}$ $\underline{69.9}$ $58.2_{\pm 0.9}$ $66.0$ $59.7_{\pm 1.2}$ $\underline{58.2}$ $59.3_{\pm 0.7}$ $62.1$ $52.2_{\pm 2.3}$ $\mathbf{67.9}$ $61.5_{\pm 1.8}$ $\underline{65.6}$ $57.3_{\pm 1.4}$ $\underline{65.0}$
+GPT-4 Turbo Smurfs $59.3_{\pm 1.4}$ $\mathbf{71.2}$ $\underline{73.3_{\pm 1.3}}$ $\mathbf{72.5}$ $\underline{67.4_{\pm 0.7}}$ $\mathbf{69.6}$ $\underline{66.7_{\pm 1.9}}$ $\mathbf{73.4}$ $\underline{55.5_{\pm 1.4}}$ $66.0$ $\underline{70.5_{\pm 0.0}}$ $\mathbf{72.1}$ $\underline{65.5_{\pm 1.1}}$ $\mathbf{70.8}$
+Table 3: The open-end tool planning task evaluation on the StableToolBench benchmark Guo et al.
+(2024).
+The most effective approach is highlighted in bold, while the second-best is underlined.
+Win rate is calculated by comparing each model with ChatGPT-ReACT.
+A win rate higher than 50% means the model performs better than ChatGPT-ReACT.
 Executor Agent
-The Executor Agent is responsible for choosing and executing the tools to solve the sub-tasks.
-The agent has access to an external tool library Qin et al.
+The Executor Agent is responsible for choosing and executing the tools to solve the sub-tasks. At each time t, the agent can invoke one tool to tackle the given task:
+$\gamma=EA.gen\_thought(p,M,\tau,h)$ (2)
+$\alpha=EA.choose\_tool(p,\gamma,\tau)$ (3)
+$\beta=EA.gen\_arguments(p,M,D[\alpha])$ (4)
+$r=EA.call\_tool(\alpha,\beta)$ (5)
+Where p is the sub-problem from Planning Agent, h is the hint from the Verifier Agent, $\tau$ is the tool list, M is local memory, $D[\alpha]$ means the tool document of tool $\alpha$.
+The agent, using the ReACT format Yao et al.
+(2022) to choose the tool and arguments, then execute the tool.
+Noticed that each inference process only uses the relevant part from the local memory and tool list to reduce the context redundancy.
+More detailed information of the Executor Agent can be found in Figure 6.
+Answer Agent
+To mitigate the performance degradation caused by lengthy contexts, we introduce the Answer Agent role, designed to extract crucial content for each step and sub-problem:
+$Answer:a=AA(q,r,M)$ (6)
+Where q is sub-problem from the Planning Agent, r is response from the Executor Agent, M is the local memory (or global memory if max retry reaches).
+As the ‘lost-in-the-middle’ theory described in section 2.1, retaining all information may not always be beneficial, particularly in cases where the solution path is challenging to discern.
+Therefore, the primary role of the Answer Agent is to succinctly summarize the generated answers and tool responses to maintain the memory efficiency.
+Verifier Agent
+The Verifier Agent serves as an early-stopping and reflection mechanism, allowing for a balance between effectiveness and efficiency:
+$h,c=VA(q,a)$ (7)
+Where q denotes the sub-problems from the Planning Agent, a denotes the answer from the answer agent, h and c denotes hint and check status respectively.
+If check status generated is 0, that means the Verifier Agent thinks the sub-problem isn’t completed, the system will add the thought and answer of this time to the local and global memory, set t=t+1 and continue the inference procedure.If check status is 1, the sub-problems is thought to be solved and the system will start to deal with the next sub-problem.
+4 Experiments
+To evaluate both the effectiveness and efficiency of the Smurfs framework, in thie section, we carried out two multi-tool planning tasks: (1) an open-ended task, StableToolBench Guo et al.
+(2024), and (2) a closed-ended task, HotpotQA Yang et al.
 (2018).
-At each steps, the agent can invoke one tool to tackle the given task.
-As outlined in Algorithm 1, the agent, using the ReACT format Yao et al.
-(2022) to choose the tool and arguments, then execute the tool.
-More detailed information of the Executor Agent can be found in Appendix B.2.
-Input: A task $q$, a hint from the verifier agent $h$, problem solving history $H$ and a set of available tools $T$.
-Output: Tool response that is useful for solving the task.
-thought $\leftarrow$ gen_thought($q$, $h$, $T$);
-tool $\leftarrow$ choose_tool($q$, thought, $T$);
-args $\leftarrow$ gen_arguments($q$, tool, $H$, tool_doc);
-response $\leftarrow$ call_tool(tool, args);
-Algorithm 1 Tool Call
-Answer Agent
-To mitigate the performance degradation caused by lengthy contexts, we introduce the ‘Answer Agent‘ role, designed to extract crucial content for each step and subtask.
-As demonstrated in the pilot study presented in Section 2.1, retaining all information may not always be beneficial, particularly in cases where the solution path is challenging to discern.
-Therefore, the primary role of the Answer Agent is to succinctly summarize the generated answers.
-Verifier Agent
-Similar to current reasoning frameworks Wei et al.
-(2022); Yao et al.
-(2022), sequential task reasoning may lead to high computational and tool resource waste within the framework Qin et al.
+In addition to these main experiments designed to assess the entire framework, we conducted an ablation studies followed by a case study to test the capabilities of each component within the multi-agent framework and investigate the underlying reasons for its effectiveness.
+4.1 Open-ended Task: StableToolBench
+StableToolBench is a tool learning benchmark derived from ToolBench Qin et al.
+(2024), encompassing multi-step tool usage tasks across over 16,000 APIs.
+The benchmark employs two metrics for evaluation: (1) Pass Rate measures the percentage of instructions successfully executed within the allocated budget.
+(2) Win Rate represents the preference selection by a ChatGPT evaluator when presented with two solution paths.
+Baselines
+Following the original paper that introduced the benchmark, we adopt ReACT (CoTWei et al.
+(2023) and DFSDT Touvron et al.
+(2023) as baseline methods for comparison.
+Additionally, we include the backbones used in the paper: gpt-3.5-turbo-0613 (GPT-3.5 Turbo) OpenAI and gpt-4-turbo-preview (GPT-4 Turbo).
+To explore the adaptability of the tool-planning methods, we also include Mistral-7B-Instruct-v0.2 (Mistral-7B) Jiang et al.
+(2023) as one of the selected backbones in our experiments.
+Settings
+To minimize the influence of varying tool APIs on experimental results, we conducted all experiments using the same API cache Guo et al.
 (2024).
-Therefore, inspired by Qin et al.
-(2024), we employ a depth-first search with an early stopping strategy to find the solution path.
-The Verifier Agent serves as an early-stopping and reflection mechanism in the intelligent multi-agent ensemble, allowing for a balance between effectiveness and efficiency.
-This mechanism not only ensures the accuracy of the generated responses but also optimizes resource use by preventing unnecessary computations.
-Moreover, if the Verifier Agent thinks the answer at this step isn’t accurate and reasonable enough, it will provide hints for the Executor Agent to get the missing information for the next steps.
-This dual role of the Verifier Agent enhances the overall performance of the framework, making it more robust and reliable for handling complex tasks.
-3.2 Subtask Solving Process
-After introducing the function of each agent, this section outlines how the agents collaborate to solve sub-tasks, as shown in Figure 3.
-Upon receiving a subtask or entering a new step, the tool list is refreshed, allowing the use of all available tools.
-The Executor Agent then calls the tools according to the task instruction.
-If a tool call fails, that tool is marked as unusable for this task at this step and the Executor Agent will try using other tools to solve the subtask at this step.
-This system also introduces a backtracking mechanism similar to DFSDT Qin et al.
-(2024) to handle the situation where errors frequently occur.
-Once the executor agent gets a correct response from a tool, the Answer Agent will refine the information, filtering out irrelevant details such as lengthy web pages and generate the answer for the subtask at this step using the local memory.
-After this, the Verifier Agent checks the answer’s accuracy.
-If it’s incorrect, the process returns to the Executor Agent with the hint from the Verifier Agent for the next step; Otherwise, the final answer is provided.
-If the Executor Agent reaches its retry limit without success, Answer Agent will review the entire process from the global memory to produce an answer.
-More details of the subtask solving process can be seen at  B
-4 Experiments
-To evaluate the effectiveness and efficiency of the Smurfs framework, we carried out a series of thorough experiments.
-In addition to the main experiment designed to assess the entire framework, we conducted an ablation study to test the capabilities of each component within the multi-agent framework.
-This section offers a detailed description of the experimental setup, methodologies employed, and key findings.
-Our goal is to showcase how the Smurfs framework, through the cooperative work of its multi-agent ensemble, effectively manages complex tasks while optimizing resource use.
-The experiments were designed with the following research objectives:
-  • To validate the capability of the entire framework in managing tool planning tasks;
-  • To independently assess the impact of each component on the overall performance, and identify the key factors influencing the multi-agent framework;
-4.1 Evaluation
-Our experiments are conducted on ToolBench Qin et al.
-(2024), which encompasses multi-step tool usage tasks across over 16,000 APIs.
-To evaluate the planning and reasoning capabilities of the LLMs, we focused our experiments on intra-category multi-tool instructions (I2) and intra-collection multi-tool instructions (I3).
-These instructions involve selecting 2-5 tools from the same category or collection and sampling up to 3 APIs from each tool to formulate the instructions.
-We employed two metrics for evaluation: (1) Pass Rate measures the percentage of instructions successfully executed within the allocated budget, evaluated by ChatGPT.
-(2) Win Rate represents the preference selection by a ChatGPT evaluator when presented with two solution paths.
-All other settings are kept consistent with those of the ToolBench benchmark.
-Models Method I2-Inst. I2-Cat. I3-Inst. Average
-Pass Win Pass Win Pass Win Pass Win
-ToolLLaMA-7B ReACT 30.5 50.8 31.5 41.8 25.0 55.0 29.0 49.2
-DFSDT 77.0 68.5 77.0 58.0 66.0 69.0 73.3 65.2
-Vicuna-7B ReACT & DFSDT 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-Smurfs (ours) 77 73 70.5 64.25 78.0 87.0 75.2 74.8
-Mistral-Instruct-7B ReACT & DFSDT 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-Smurfs (ours) 77.5 80.0 79.5 79.2 79.0 94.0 78.7 84.4
-GPT4 ReACT 67.0 65.8 72.0 60.3 47.0 78.0 62.0 68.0
-DFSDT 79.5 73.3 77.5 63.3 71.0 84.0 76.0 73.5
-Smurfs (ours) 71.0 77.5 72.0 77.0 64.0 89.5 69.0 81.3
-Table 3: ToolBench evaluation, with some results derived from Qin et al. (2024); Yuan et al. (2024). The most effective approach is highlighted in bold, while the second is underlined.
-4.2 Baselines
-To investigate the varying impacts of the agent framework on models with different capabilities, we categorize our baseline into three groups.
-The first group consists of models that are fine-tuned based on the tool dataset, represented by ToolLLaMA Qin et al.
+For a fair comparison among the candidate methods and to reduce variability, each model was executed once and evaluated three times, with results averaged.
+Other settings follow those specified in the original benchmark paper.
+Results
+Table 3 displays the results on StableToolBench.
+For the untrained LLM, Mistral-7B, existing agent frameworks did not improve its performance in tool planning tasks; Mistral-7B failed these tasks when integrated with the ReACT and DFSDT frameworks ***Experiment results show that Mistral-7B failed to correctly execute the ‘finish’ action during inference, resulting in invalid responses..
+However, Smurfs exhibited exceptional performance: when combined with Mistral-7B, Smurfs achieved competitive scores among the baselines.
+Through its task decomposition mechanism, Smurfs transforms long-context tasks into simpler ones, enabling the untrained model to effectively utilize external tools for managing complex tasks.
+Regarding closed-source models, specifically GPT4 in these experiments, Smurfs also demonstrated outstanding performance on the benchmark compared to other agent frameworks and achieved state-of-the-art results on the benchmark.
+Its high success rate suggests that Smurfs is more effective at finding optimal solution paths compared to ChatGPT.
+Further Analysis
+We conducted a detailed analysis of the token costs associated with each tool planning method for the tasks, a critical evaluation aspect for multihop reasoning tasks.
+As shown in Table 1 (detailed in Appendix E), the average token costs per question and API request are evaluated for ReACT, DFSDT, and Smurfs on StableToolBench.
+The analysis reveals that DFSDT generally requires about 20,000 tokens per question, encompassing both prompt and completion tokens.
+This is nearly three times the token cost compared to ReACT and twice as much as Smurfs.
+Despite this higher token cost, DFSDT does not demonstrate commensurate effectiveness improvements over other methods.
+These findings underscore the cost-efficiency of the proposed MAS framework, Smurfs, which not only reduces token expenditure in solving multihop planning tasks but also delivers outstanding performance in evaluations.
+[…]
+Easy Medium Hard All
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+4.2 Closed-ended Task: HotpotQA
+Compared to open-ended tasks, closed-ended tasks provide a more stable and robust evaluation.
+To this end, we evaluate the methods on HotpotQA Yang et al.
+(2018) in addition to StableToolBench.
+HotpotQA is a multi-hop QA task that is challenging due to the requirement for rich background knowledge, with answers typically being short entities or yes/no responses.
+Baselines
+The compared baselines include CoT Wei et al.
+(2023), ReActYao et al.
+(2022), ChameleonLu et al.
+(2023), Reflexion Shinn et al.
+(2023), BOLAA Liu et al.
+(2023), ReWOO Xu et al.
+(2023), FireAct Chen et al.
+(2023a), AutoActQiao et al.
 (2024).
-The second group encompasses untrained general language models such as Vicuna Chiang et al.
-(2023), and Mistral-Instruct-7B Jiang et al.
-(2023).
-The third category represents the closed-source model, embodied by GPT4.
-We subsequently contrast our approach with two agent frameworks, ReACT Yao et al.
-(2022) and DFSDT Qin et al.
-(2024), both of which are utilized for multi-step reasoning and model invocation.
-Notably, all methods employ the ground truth toolset for tool selection, thereby eliminating the influence of the tool retriever.
-4.3 Main Experiments
-Table 3 displays the results of the comprehensive evaluation of our proposed framework on ToolBench.
-For the untrained LLMs, it is clear that existing agent frameworks do not improve their performance in tool planning tasks; Vicuna, and Mistral-Instruct-7B all failed at the given tasks with the ReACT and DFSDT frameworks.
-However, Smurfs exhibits exceptional performance: Mistral combined with Smurfs achieves the highest score among the baselines.
-Through its task decomposition mechanism, Smurfs transforms a long-context task into several simpler tasks, enabling the untrained model to effectively utilize external tools to manage complex tasks.
-Regarding the closed-source models, specifically GPT4 in these experiments, Smurfs also demonstrates competitive performance on the benchmark compared to other agent frameworks.
-A high win rate suggests that Smurfs is more adept at finding better solution paths than ChatGPT, a success that can likely be credited to the verifier agent.
-4.4 Ablation Study
-4.4.1 Importance of each component in MAS
-Average
-Pass Win
-Mistral 0.0 0.0
-Mistral with Smurfs 79.3 86.6
-   w/o Verifier Agent $74.5_{\downarrow 6.0\%}$ $83.8_{\downarrow 3.2\%}$
-   w/o Answer Agent $73.3_{\downarrow 7.6\%}$ $81.7_{\downarrow 5.7\%}$
-   w/o Planning Agent $64.0_{\downarrow 19.3\%}$ $82.9_{\downarrow 4.3\%}$
-Table 4: Ablation study on ToolBench I2-Cat., I3-Inst. with Mistral-Instruct-7B.
+Settings and Metrics
+Following the settings in  Qiao et al.
+(2024), we use open-source Llama-2 models Touvron et al.
+(2023) and Mistral-7B Jiang et al.
+(2023) as the backbones of each agent to evaluate the performance of Smurfs.
+The evaluation metrics is $\texttt{reward}\in[0,1]$, defined as the F1 score grading between the prediction and ground-truth answer.
+For more details about the experiment, see Appendix C.
+Results
+Smurfs, as an untrained MAS system, not only comprehensively outperforms untrained agents but also achieves and even surpasses the accuracy of trained agents across most backbone models.
+This sufficiently demonstrates that the mechanism of smurfs ensures strong generalization capabilities while maintaining high effectiveness.
+Observations indicate that the performance of LLama-2-13b-chat on smurfs-related tasks is suboptimal, likely due to its limited capabilities in tool arguments generation.
+Specifically, the primary issue identified is that, when the Executor agent successfully selects relevant tool, it tends to produce hallucination arguments that can’t be used by the tools.
+This indicates that LLama-2-13b-chat may need further training for usage of tools.
+The experimental results may substantiate this viewpoint, demonstrating that the untrained methods of llama-2-13b-chat generally exhibit significantly lower accuracy compared to the trained methods.
+Nevertheless, Smurfs achieves the second highest accuracy among the untrained methods, only slightly behind reflexion, which still attests to Smurfs’ capability.
+I3-Inst.
+[…]
+GPT-3.5 Turbo with Smurfs $60.1_{\pm 1.0}$ 57.4
+   w/o Answer Agent $57.4_{\pm 2.9}$ $49.2$
+   w/o Verifier Agent $54.1_{\pm 2.7}$ $42.6$
+   w/o Planning Agent $35.5_{\pm 3.3}$ $42.6$
+GPT-4 Turbo with Smurfs $70.5_{\pm 1.0}$ 72.1
+   w/o Answer Agent $82.2_{\pm 2.5}$ $72.1$
+   w/o Verifier Agent $79.2_{\pm 0.8}$ $63.9$
+   w/o Planning Agent $71.9_{\pm 2.8}$ $63.9$
+Table 5: Ablation study on StableToolBench I3-Inst subset to investigate the importance of each component within the framework.
+Figure 3: The illustration of how GPT4-Smurfs and GPT4-DFSDT solve long context problem. The two sub-questions and their corresponding answers are marked in two colors.
+5 Ablation Study
+5.1 Importance of each component in MAS
 We performed an ablation study to investigate the impact of each agent in our framework.
 We removed each agent individually, except for the indispensable Executor Agent, and compared the results to the complete framework.
-Table 4 shows that the Planning Agent is the most crucial component, followed by the Answer Agent and the Verifier Agent.
+Table 5 shows that the Planning Agent is the most crucial component, followed by the Verifier Agent, with the Answer Agent being the least important.
 (1) Verifier Agent Removal: Without verification, the framework uses a general depth-first search, leading to increased computational demand and more tool invocations.
-(2) Answer Agent Removal: Removing this agent means the Executor Agent’s answer won’t be summarized, risking the ’lost-in-the-middle’ problem due to lengthy tool responses.
-(3) Planning Agent Removal: Without this agent, the global path searching strategy is affected.
+(2) Answer Agent Removal: Removing this agent means the Executor Agent’s answers won’t be summarized, risking the ’lost-in-the-middle’ problem due to lengthy tool responses.
+As shown in the results, a more intelligent model, GPT-4 Turbo, can mitigate the negative impact of the Answer Agent’s removal.
+We believe this is because the more powerful model can leverage more information effectively.
+(3) Planning Agent Removal: Removing the Planning Agent affects the global path-searching strategy.
 Models with Smurfs may show reduced performance without preliminary planning, as seen in current frameworks like ReACT and DFSDT.
-4.4.2 Effect of number of agents
-Models # of I2-Cat. I3-Inst. Average
-Agents Pass Win Pass Win Pass Win
-Vicuna 1 49.5 53.0 60.0 86.0 54.8 69.5
-2 71.0 66.0 75.0 83.0 73.0 74.5
-3 70.5 64.3 78.0 87.0 74.3 75.6
-Mistral 1 65.5 70.5 56.0 94.0 60.8 82.3
-2 76.0 75.5 73.0 92.0 74.5 83.8
-3 79.5 79.2 79.0 94.0 79.3 86.6
-GPT4 1 60.5 71.8 57.0 89.0 58.8 80.4
-2 69.5 77.8 59.0 94.0 64.3 85.9
-3 72.0 77.0 64.0 89.5 68.0 83.3
-Table 5: The impact of MAS complexity on the performance of different models. More agents don’t always bring more benefits.
-Figure 4: The illustration of how GPT4-Smurfs and GPT4-DFSDT solve long context problem. The two sub-questions and their corresponding answers are marked in two colors.
-In addition to evaluating the importance of each component, another relevant question arises: Can a fixed agent system effectively accommodate every model?
-Our hypothesis suggests that larger language models possess comprehensive intelligence, and employing complex agent systems may potentially hinder their performance compared to simpler ones.
-As shown in Table 5, we conducted an ablation study on Vicuna, Mistral-Instruct-7B, and GPT4, varying the number of agents ***Agents are incrementally added to the system based on their importance..
-The findings indicate that as the number of agents in a MAS increases, leading to increased complexity of the MAS, the system’s performance enhancement does not scale linearly.
-In certain instances, GPT4 with a 2-agent MAS outperforms the one with a 3-agent MAS, suggesting that larger language models do not consistently benefit from more complex Multi-Agent Systems.
-Smaller language models can benefit more from the agent system, as seen with Vicuna increasing from 69.5 to 75.6.
+The results demonstrate that the impact of removing the Planning Agent is significant, as it directly influences the multihop reasoning ability of the MAS.
 5.2 Case Study
-As shown in Figure 4, even though GPT4-DFSDT and GPT4-Smurfs use the same tool calls to solve the problem, GPT4-DFSDT only answers the first sub-question correctly while GPT4-Smurfs answers both sub-questions accurately.
+As shown in Figure 3, although GPT4-DFSDT and GPT4-Smurfs use the same tool calls to solve the problem, GPT4-DFSDT only answers the first sub-question correctly while GPT4-Smurfs answers both sub-questions accurately.
 In the process of addressing the second sub-question, it is notable that the tool response only mentions titles of film and television products related to "Star Wars", without addressing OTT platforms.
@@
 Conversely, the context-efficient Smurfs framework demonstrates a reduced susceptibility to irrelevant information, thereby generating more accurate answers.
-5 Related Work
-In this section, we review the literature related to multi-agent collaboration, tool-augmented language models, and complex task planning and reasoning.
-Each subsection provides an overview of the topic, discusses recent advancements, identifies challenges, and suggests avenues for future research.
-5.1 Multi-Agents Collaboration
-Multi-agent systems have garnered significant attention due to their applicability in various domains such as robotics, economics, and computer networking.
-One of the key challenges in multi-agent systems is achieving effective collaboration among autonomous agents.
-Recent studies have focused on understanding collaborative behaviors and developing coordination strategies to enable agents to work together towards common goals.
-For example, research in reinforcement learning has explored techniques for emergent coordination, where agents learn to collaborate through interaction with the environment and other agents.
-Du et al.
-(2023); Liang et al.
-(2023) carry out multi-agent interaction in the form of debate, which improves mathematical and strategic reasoning tasks.
-Li et al.
-(2023); Wang et al.
-(2023) uses role-playing to conduct the interaction between multi-agents.
-5.2 Tool-Augmented Language Models
-Language models augmented with external knowledge sources, often termed as tool-augmented language models, have demonstrated promising outcomes in a variety of natural language processing tasks.
-These models utilize external knowledge bases, ontologies, or pre-trained models to enhance their contextual understanding and boost performance in tasks such as text generation, summarization, and question answering.
-Recent advancements in this field include techniques for integrating structured knowledge into language models, such as graph-based representations or semantic parsing.
-For instance, Qin et al.
-(2024) introduces ToolBench, an instruction-tuning dataset for tool usage, along with a fine-tuned tool-oriented model, ToolLLaMA.
-Another significant contribution is Gorilla Patil et al.
-(2023), which excels at writing API calls and also introduces a benchmark for evaluating LLMs with tools.
-5.3 Complex task planning and reasoning
-Problem decomposition (Zhou et al.
-(2023); Drozdov et al.
-(2022); Khot et al.
-(2022); Press et al.
-(2022)) is a popular paradigm used in the LLM and leads to good performance in challenging reasoning tasks.
-This divide-and-conquer strategy provides sufficient explanation of how the model works.
-The current trend is that when model size reaches some limit, people start looking for a way to make the most of the limited model.
-Chain-of-thought prompting Wei et al.
-(2022) is one of the strategy that lead models to divide the tasks into a chain, and solve the task step by steps.
-Chen et al.
-(2022) propose a program of thoughts structure to enhance the reasoning ability.
-Yao et al.
-(2024) simulate thought as tree structure to solve task.
-Besta et al.
-(2024) using graph to simulation.
 6 Conclusion
-In this study, we present a novel MAS framework, Smurfs, tailored to enhance the planning and reasoning capabilities of LLMs in handling complex tasks that involve lengthy contexts and tools.
-We conducted experiments on the multi-step tool usage benchmark, ToolBench, and the results demonstrated the overall effectiveness of the Smurfs framework compared to the baseline models.
-Ablation studies are carried out to investigate and compare the significance of different components within the MAS framework.
-The findings revealed that preliminary planning was the most crucial element.
-Content summarizing also played a key role in mitigating the ’lost-in-the-middle’ issue often encountered in long-context multi-step reasoning scenarios.
-While verification was not as influential on effectiveness, it proved valuable in enhancing computational efficiency by identifying the optimal solution path for complex task resolution.
-By dissecting and comparing the different aspects of the MAS framework, we aim to offer insights that could inspire advancements in the applicability and accuracy of LLMs.
+In this study, we present a novel MAS framework, ‘Smurfs’, tailored to enhance the planning and reasoning capabilities of LLMs in handling complex tasks that involve lengthy contexts and tools.
+We conduct experiments on the multi-step tool usage benchmark, StableToolBench and HotpotQA, and the results demonstrate the overall effectiveness and efficiency of the Smurfs framework compared to baseline methods.
 In conclusion, this research contributes to the expanding field of study focused on enhancing LLM capabilities, particularly for multi-step tool usage tasks.
 It emphasizes the importance of task decomposition, preliminary planning, and efficient verification for improving task execution performance.
-We are confident that the knowledge gained from this study will lay the groundwork for the development of more sophisticated and efficient LLM frameworks in the future.
+For future work, we believe incorporating more dedicated and specific roles within the system may further enhance effectiveness and efficiency, based on the ‘Smurfs principle’: synergistic collaboration among specialized agents can overcome the limitations faced by individual LLMs.
 7 Limitations
-Generalization Ability:
-While our empirical investigation has demonstrated promising results for the Mistral-7b-instruct model, additional benchmark evaluations may be needed to validate the generalization ability of the proposed Smurfs framework across various LLM architectures and tasks.
 Model Size Constraints:
-Due to device limitations and computational constraints, our experiments primarily focused on the 7B models.
-Further evaluations with larger and smaller LLMs are required to assess the impact of the Smurfs framework on models of different sizes.
-Computational Efficiency:
-Although preliminary findings suggest that the Smurfs framework can enhance computational efficiency through efficient verification, a more detailed analysis is needed to quantify the computational overhead introduced by the multi-agent architecture.
+Due to computational constraints, our experiments did not include larger and more diverse types of LLMs.
+Agent Component Scale-Up:
+Although we selected the most common and intuitive agent roles for the proposed MAS, there are many possibilities for researchers to explore.
+Investigating more well-designed agent roles may help improve the effectiveness of the agent system, and developing automated methods to identify these roles could facilitate effective scaling.
 Acknowledging these limitations, future research should aim to address these gaps to provide a more comprehensive understanding of the Smurfs framework’s capabilities and potential areas for improvement.
-Appendix A Pilot Study Settings
-Figure 5: Demonstration of the memory of the Smurfs framework.
-A.1 Pilot Study 1
-We conducted the first pilot study using the ToolBench benchmark Qin et al.
-(2024), which comprises over 16,000 APIs focused on tool usage and planning.
-We divided the benchmark into two subsets: "ToolBench-long," containing samples requiring more than three steps for ChatGPT-DFSDT, and "ToolBench-all," which includes all samples from ToolBench.
-On average, we consider the "ToolBench-long" subset to be more complex than the "ToolBench-all" set.
-We maintained all other settings consistent with the original source settings.
-A.2 Pilot Study 2
-The second pilot study is conducted using the BMTools Qin et al.
-(2023) test datasets.
-We augmented each sample by adding random, irrelevant test samples as prefix prompts to increase their input length.
-We used tool selection accuracy as the metric to evaluate the LLMs’ model selection capabilities.
-The experiment uses Mistral-7B-Instruct-v0.1 Jiang et al.
-(2023) and Qwen1.5-7B-Chat Bai et al.
-(2023) for evaluation.
+Appendix A Details of DFSDT
+See Figure 5.
+Figure 5: A comparison of DFSDT and conventional CoT or ReACT during model reasoning (left) Qin et al. (2024). We show part of the solution path annotation process using GPT-4 (right).
 Appendix B Details of the Smurfs
+See Figure 6 for executor working process and Figure 4 for memory and tool library of Smurfs.
 Figure 6: Details of the executor agent working process
-B.1 Memory management of Smurfs
-When each agent performs a task, the prompt information they receive varies.
-We refer to this as the memory management system of Smurfs.
-As illustrated in Figure 5, information is primarily divided into four categories.
-The first category is local memory, consisting of the thought and answer from the previous steps.
-When backtracking happens, the thought and answer of the backtracking step will be popped out from the local memory.
-The second category is global memory, which stores all the previous steps history including those that have been backtracked.
-The third category is the tool list, similar to other tool-using frameworks, storing brief information about all accessible tools.
-Lastly, the tool document provides detailed usage information like the parameters infomation for each tool listed.
-As illustrated in Figure 3, each agent has its own memory context.
-The Executor Agent’s memory will be discussed in the next section.
-The Answer Agent takes in the subtask, the tool response and the local memory to generate the answers to the subtask using all the local memory so far.
-Then the thought and answer of this step will be added to the local memory and the global memory.
-The Verifier Agent takes in the subtask, the answer from the Answer Agent and return the status of the subtask.
-If the status is solved, then the system will return the subtask’s answer and process to the next subtask; If the status is pending, the Verifier Agent will give hint to the Executor Agent to process to the next step.
-When the max steps have been reached for a subtask, the answer agent will generate the answer for the subtask using the global memory and process to the next subtask.
-B.2 Executor Agent Workflow
-Next, let’s illustrate how memory is accessed during the Executor Agent’s working process, as shown in Figure 6.
-When a subtask arises, the Executor Agent first receives the hint (task instruction) from the Verifier Agent, the subtask, the available tool list to generate the thought(a strategy for task execution) at this step.
-It then takes in the subtask, the thought and the tool list to generate an action with a specific action name.
-Subsequently, it takes in the subtask, the thought, the tool document of the chosen action and the local memory to generate the input of the chosen action.
-B.3 Backtracking Mechanism
-Smurfs uses a backtracking mechanism similar to DFSDT  Qin et al.
+Appendix C Experiment Settings for Hotpot QA
+Following settings in Qiao et al.
+(2024), which is randomly select 300 dev questions divided into three levels for evaluation, with 100 questions in each level.
+For tool library that can be used in HotpotQA, see Table 6
+Name Definition Usage
+  BingSearch BingSearch engine can search for rich knowledge on the internet based on keywords, which can compensate for knowledge fallacy and knowledge outdated.
+BingSearch[query], which searches the exact detailed query on the Internet and returns the relevant information to the query.
+Be specific and precise with your query to increase the chances of getting relevant results.
+For example, Bingsearch[popular dog breeds in the United States]
+Retrieve Retrieve additional background knowledge crucial for tackling complex problems.
+It is especially beneficial for specialized domains like science and mathematics, providing context for the task Retrieve[entity], which retrieves the exact entity on Wikipedia and returns the first paragraph if it exists.
+If not, it will return some similar entities to retrieve.
+For example, Retrieve[Milhouse]
+Lookup A Lookup Tool returns the next sentence containing the target string in the page from the search tool, simulating Ctrl+F functionality on the browser.
+Lookup[keyword], which returns the next sentence containing the keyword in the last passage successfully found by Retrieve or BingSearch.
+For example, Lookup[river].
+Table 6: Tool library for HotpotQA.
+Appendix D Prompts for multi-agent implementation
+Prompts used by each agent and their example outputs are shown in Figure  7 to  13.
+Appendix E Token Cost on StableToolBench Evaluation
+We analyzed the token cost for the StableToolBench experiments.
+As shown in Table 7, the total token cost for each subtask within the StableToolBench is compared across three candidate tool-planning methods.
+The data demonstrates that, across all tasks from easy to hard, DFSDT consistently incurs high token costs, while the other two methods maintain relatively low token costs.
+This verifies the context-efficiency of the proposed method.
+Backbone Method StableToolBench
+I1-Inst. I1-Cat. I1-Tool. I2-Cat. I2-Inst. I3-Inst. Average
+Total Avg. Total Avg. Total Avg. Total Avg. Total Avg. Total Avg. Total Avg.
+GPT-3.5 Turbo ReACT 1,010,304 6,198 824,676 5,390 1,010,514 6,396 900,855 7,265 824,510 7,778 461,121 7,559 838,663 6,764
+GPT-3.5 Turbo DFSDT 3,303,062 20,264 2,745,667 17,945 3,152,532 19,953 2,560,297 20,648 3,098,365 29,230 1,390,787 22,800 2,708,452 21,807
+GPT-3.5 Turbo Smurfs 1,090,404 7,127 1,917,348 11,763 1,464,535 9,269 957,088 7,638 1,096,162 10,341 632,084 10,362 1,191,270 9,417
+Table 7: Token costs for various candidate tool-planning methods on the StableToolBench benchmark Guo et al.
 (2024).
-As illustrated in Figure 3, when tool responses error, the system will delete the tool from the tool list and retry this step.
-However, when all tools have reported error or the model thinks the available tools can’t solve the subtask at this step, the system will pop out the latest local memory and the latest step, delete the tool used by the popped out step from the tool list of the current step and retry the the other solution path.
-Appendix C Prompts for multi-agent implementation
-Prompts used by each agent and their example outputs are shown in Figure  7 to  13.
+‘Total’ indicates the total number of tokens used to complete each subtask, including both prompt and completion tokens.
+‘Avg.’ represents the average number of tokens used per question within the subtasks.
+Higher token counts imply greater costs for solving the same task.
 Planning Agent Prompt: You need to decompose a complex user’s question into some simple sub-tasks and let the model execute it step by step.

### v4

@@
-Smurfs: Leveraging Multiple Proficiency Agents with Context-Efficiency for Tool Planning
+Smurfs: Multi-Agent System using Context-Efficient DFSDT for Tool Planning
 Abstract
-The emergence of large language models (LLMs) has opened up unprecedented possibilities for automating complex tasks that are often comparable to human performance.
-Despite their capabilities, LLMs still encounter difficulties in completing tasks that require high levels of accuracy and complexity due to their inherent limitations in handling multifaceted problems single-handedly.
-This paper introduces ‘Smurfs’, a cutting-edge multi-agent framework designed to revolutionize the application of LLMs.
-By seamlessly transforming a conventional LLM into a synergistic multi-agent ensemble, Smurfs can enhance the model’s ability to solve complex tasks at no additional cost.
-This is achieved through innovative prompting strategies that allocate distinct roles within the model, thereby facilitating collaboration among specialized agents and forming an intelligent multi-agent system.
-Our empirical investigation on both open-ended task of StableToolBench and closed-ended task on HotpotQA showcases Smurfs’ superior capability in intricate tool utilization scenarios.
-Notably, Smurfs outmatches all the baseline methods in both experiments, setting new state-of-the-art performance.
-Furthermore, through comprehensive ablation studies, we dissect the contribution of the core components of the multi-agent framework to its overall efficacy.
-This not only verifies the effectiveness of the framework, but also sets a route for future exploration of multi-agent LLM systems.
-Our code is available at https://github.com/FreedomIntelligence/Smurfs/.
-Smurfs: Leveraging Multiple Proficiency Agents with Context-Efficiency for Tool Planning
-Junzhi Chen [^1] These authors contributed equally to this work., Juhao Liang [^1] These authors contributed equally to this work., Benyou Wang🖂 Shenzhen Research Institute of Big Data The Chinese University of Hong Kong, Shenzhen wangbenyou@cuhk.edu.cn
+Teaching large language models (LLMs) to use tools for solving complex problems can grant them human-like reasoning abilities.
+ReAct and its variants are popular frameworks for tool use in both single-agent and multi-agent systems.
+To address issues like error propagation and limited exploration in ReAct, the Deep First Search Decision Tree (DFSDT) was proposed, but it faces challenges such as rollback instability, redundant context, and premature termination in single-agent settings.
+We introduce "Smurfs," a novel multi-agent system (MAS) that enhances DFSDT with a modular, context-efficient, and training-free design.
+Smurfs surpasses baseline methods in both the open-ended StableToolBench and the closed-ended HotpotQA tasks, reducing token usage by 60.9% compared to DFSDT and enabling Mistral-7b to perform on par with GPT-4-DFSDT.
+Extensive ablation studies confirm the effectiveness of Smurfs’ core components, offering valuable insights for the construction and interpretation of MAS, and paving the way for future exploration.
+We release the code at https://github.com/FreedomIntelligence/Smurfs.
+Smurfs: Multi-Agent System using Context-Efficient DFSDT for Tool Planning
+Junzhi Chen [^1] These authors contributed equally to this work., Juhao Liang [^1] These authors contributed equally to this work., Benyou Wang🖂 The Chinese University of Hong Kong, Shenzhen Shenzhen Research Institute of Big Data wangbenyou@cuhk.edu.cn
 footnotetext: 🖂Corresponding author.
@@
 1 Introduction
-Tool manipulation has traditionally been seen as a distinctive human characteristic, dating back approximately 2.5 million years Oakley and Museum (1972); Ambrose (2001).
-For large language models (LLMs), access to external tools can equip them with broader capabilities beyond their fixed language modeling knowledge.
-For example, the search engine API empowers ChatGPT to access real-time information Zhao et al.
-(2023).
-However, LLMs still encounter several challenges when using multiple tools to solve tasks.
-These challenges include effective solution planning and adaptability to new tools.
-Hao et al.
-(2024); Guu et al.
-(2020); Qin et al.
-(2024).
-This paper addresses the critical research problem of enhancing the problem-solving capabilities of LLMs through the adoption of a plug-and-play multi-agent system (MAS) framework Dorri et al.
-(2018); Van der Hoek and Wooldridge (2008).
-We posit that a MAS approach can significantly augment the efficacy of LLMs in handling tasks that require a high degree of precision, adaptability, and comprehensive knowledge integration.
+The ability to manipulate tools for complex tasks has long been considered a distinctive characteristic of humans Oakley and Museum (1972); Ambrose (2001).
+Can we extend this ability to today’s large language models (LLMs), enabling them to utilize multiple tools to perform complex tasks beyond their inherent capabilities?
+If LLMs can use external tools to access knowledge or execute tasks beyond their fixed language modeling capabilities, we can shift the focus of LLM training towards enhancing their reasoning and tool-use skills.
+This shift would allow tools to supplement what LLMs should know or execute, thereby improving the parameter efficiency of the LLMs.
 Pass Rate $\uparrow$ (%) Win Rate $\uparrow$ (%) # of Tokens per request $\downarrow$ # of Tokens per query $\downarrow$
@@
 The results are averaged over the subtasks within StableToolBench.
-To this end, we introduce ‘Smurfs’ an innovative MAS framework inspired by the collaborative and versatile nature of its namesake cartoon characters.
-The proposed framework is based on the principle: synergistic collaboration among specialized agents can overcome the limitations faced by individual LLMs.
-Each agent within the Smurfs framework is designed to perform specific sub-tasks, facilitating a more nuanced and effective approach to complex problem-solving.
-Our research delves into the architectural design, coordination mechanisms, and the operational dynamics of integrating specialized agents into a cohesive system.
-The effectiveness of Smurfs is validated through both open-ended and closed-ended tool planning benchmark experiments Guo et al.
+In previous multi-agent systems for tool planning, methods like Chain-of-Thought (CoT) Wei et al.
+(2023), ReAct Yao et al.
+(2022), and the more advanced DFSDT Qin et al.
+(2024) have been proposed to enhance LLMs’ ability to handle complex multi-step tasks.
+However, these approaches face notable limitations.
+ReAct has trouble in handling error propagation and limited exploration.
+DFSDT incorporates a rollback mechanism and depth-first search approach to address limitations of ReAct, but it suffers from instability when the base model struggles with long-context reasoning.
+It also introduces inefficiencies due to redundant context handling and risks premature termination when solving multi-step problems.
+These challenges highlight the need for further innovation to better manage context and reasoning complexity in multi-tool planning systems.
+In this paper, we introduce ‘Smurfs’, an innovative multi-agent system (MAS) framework inspired by the collaborative and versatile nature of its namesake cartoon characters.
+The proposed framework leverages enhanced DFSDT to perform complex tool planning tasks without the need for additional training.
+The effectiveness of Smurfs is demonstrated through both open-ended and closed-ended tool planning benchmark experiments Guo et al.
 (2024); Yang et al.
-(2018), where the proposed MAS system consistently outperform baseline methods on both benchmarks.
-An ablation study followed by a case study further investigates the underlying reasons for this effectiveness.
-These results not only establish a new state-of-the-art in the field but also offer concrete evidence of the multi-agent approach’s efficacy in enhancing LLM capabilities.
+(2018), where it consistently outperforms baseline methods.
+An ablation study, followed by a case study, further investigates the reasons behind this effectiveness.
+These results establish a new state-of-the-art in the field and provide concrete evidence of the advantages of a multi-agent approach in enhancing LLM capabilities.
 The contributions of this paper can be summarized as follows:
   1. 1.
-    We introduce a novel plug-and-play MAS framework to enhance the tool planning capabilities of LLMs.
-Experiments demonstrate the effectiveness of this approach, which is also more cost-efficient compared to existing tool planning methods.
+    We introduce a highly modular, context-efficient, and training-free MAS framework that utilizes an enhanced DFSDT to improve the tool planning capabilities of LLMs.
+Experiments demonstrate the effectiveness of this approach, which also proves to be more cost-efficient compared to existing methods.
   2. 2.
-    Ablation studies further reveal the underlying reasons for the effectiveness of the MAS framework, providing valuable insights for future research.
-2 Motivation
-2.1 Multi-Tool Planning
+    Through ablation studies, we uncover the underlying factors contributing to the effectiveness of the MAS framework, offering valuable insights for future research.
+2 Related Work
 To augment LLMs to do multi-tool planning for solving complex problems, previous work has seen numerous attempts.
 Chain-of-Thought  Wei et al.
-(2023) was the first to propose the method of thought and answer chain reasoning.
+(2023) is the first to propose the method of thought and answer chain reasoning.
 ReAct Yao et al.
-(2022) further introduced the thought-action-observation format for tool chain reasoning, leading to the development of various multi-tool planning methods Chen et al.
+(2022) further introduce the thought-action-observation format for tool chain reasoning, leading to the development of various multi-tool planning methods Chen et al.
 (2023a); Xu et al.
@@
 (2023).
-The latest work, DFSDT Qin et al.
-(2024), was proposed to address the inherent limitations of CoT and ReACT: error propagation and limited exploration.
-Deep First Search Decision Tree, denoted as DFSDT, is powerful in addressing multi-tool planning problems.
-Its core concept involves employing a depth-first search (DFS) approach for multi-tool planning (for more details, see Appendix A).
-When a tool fails or is deemed inadequate for solving the current problem, DFSDT backtracks to the previous solution state and attempts to resolve the issue using a different tool.
-However, several limitations were identified with the mechanism of DFSDT: (1) instability of the rollback mechanism, (2) redundant context, and (3) premature termination.
-The following sections will introduce these limitations in detail.
-2.1.1 Instability of the Rollback Mechanism
+The latest work, Deep First Search Decision Tree (DFSDT) Qin et al.
+(2024), is proposed to address the inherent limitations of CoT and ReACT: (1) Error Propagation: Error occurs at early stage of planning will result in wrong answer in the end, but it can only be identified until reaching the end of the planning chain (2) Limited Exploration: Single solution chain can’t explore the planning space completely.
+DFSDT is powerful in addressing multi-tool planning problems.
+Its core concept involves employing a depth-first search (DFS) approach for multi-tool planning and backtracks whenever an LLM think the solving process has entered a wrong state (for more details, see Appendix B).
+When a tool fails or is deemed inadequate for solving the current problem, DFSDT backtracks to the previous solution state and attempts to resolve the issue using another solution plan.
+However, three limitations are identified with the mechanism of DFSDT:
+Limit I: Instability of the Rollback Mechanism
 The rollback mechanism in DFSDT is determined by the model.
@@
 retry the same error tools or roll back too far.
-2.1.2 Redundant Context
+Limit II: Redundant Context
 In the process of planning with DFSDT, each tool plan is generated using the entire conversation history (including all the thoughts, actions, action inputs and tool responses) as context.
@@
 (2023), which is known as the ‘lost-in-the-middle’ problem.
-2.1.3 Premature Termination
+Limit III: Premature Termination
 The termination mechanism set by DFSDT involves adding a termination tool to the model’s selectable toolkit.
@@
 Instead, the model focuses on whether the current sub-problem requires termination, leading the mechanism to terminate after resolving the sub-problem.
-2.2 Multi Agent System
+In conclusion, DFSDT relies highly on the base model’s reasoning ability, especially long context reasoning ability to make roll back decision, termination decision and tool choice decision at the same time, which is a very difficult task even with the most powerful LLM like GPT4 Yuan et al.
+(2024).
+Multi-Agent System
+To address the limitations inherent in DFSDT and to further enhance LLM’s multi-tool planning capabilities, MAS has emerged as a natural solution.
+Inspired by human social division of labor and cooperation, MAS aim to enable AI agents to accomplish more complex tasks through the division of labor and collaboration.
+By decomposing the task of DFSDT to multiple agents and giving them only the information they need, we can enable LLMs to use DFSDT more effectively and more efficiently.
 Method Multi-Agent Training Generality Reflection Planning
@@
 Smurfs (Ours) Iterative and Global
-Table 2: Comparison of related works.
-To address the limitations inherent in DFSDT and to further enhance LLM’s multi-tool planning capabilities, multi-agent system (MAS) has emerged as a natural solution.
-Inspired by human social division of labor and cooperation, MAS aim to enable AI agents to accomplish more complex tasks more effectively and efficiently through the division of labor and collaboration.
-Previous works Song et al.
+Table 2: Comparison of tool use systems.
+Previous work on multi-agent system mainly focus on coding and society simulation area Hong et al.
+(2023); Qian et al.
+(2024); Park et al.
+(2023); Li et al.
+(2023); Wu et al.
+(2023).
+For tool-use scenario, most multi-agent systems still use the ReAct style reasoning Qiao et al.
+(2024); Chen et al.
+(2023b); Yuan et al.
 (2024); Liu et al.
-(2023); Chen et al.
-(2023b); Yin et al.
-(2024); Qiao et al.
-(2024) has leveraged MAS to achieve this goal.
-Table 2 shows the difference between them.
-Based on those works, we further design the MAS named Smurfs to address issues with DFSDT.
-By dividing tasks among different agents, each agent can focus on a specific part of the DFSDT task, accessing only the necessary history as context during task execution, which effectively addresses the issue of redundant context.
+(2023); Song et al.
+(2023); Yin et al.
+(2024); Xu et al.
+(2023), only using multi-agent discussion and revision to increase reasoning quality, which still inherits the limitation of ReAct.
+Therefore, this paper aims to construct a novel MAS framework to address the aforementioned limitations.
+Table 2 shows the difference between different tool-use systems.
+More detailed discussion can be seen in Appendix A.
+Token Compression
+Token compression refers to compressing tokens fed into LLMs while preserving inference performance, thus reducing computational overhead and mitigating the constraints imposed by long context limits.
+Previous works Mu et al.
+(2024); Fu et al.
+(2024) have explored token compression techniques, with a focus on compressing token embeddings.
+In contrast, Smurfs achieves context compression by filtering the input context for each tool planning process.
+Table 6 shows the detailed difference between different compression techniques.
+In the future, these compression methods can be used together in multi-tool planning scenarios to achieve more efficient token compression.
+Each agent’s system prompt could be compressed using gist tokens Mu et al.
+(2024), tool descriptions could be managed with CAMPHOR’s compression approach Fu et al.
+(2024), and the context input by each agent could be compressed using Smurfs.
+Figure 2: Details of the subtask-solving process of the Smurfs framework. The dotted line represents that the agent can see the memory and the full line stands for operation.
+3 Smurfs: MAS with Context Efficient DFSDT
+The Smurfs, the beloved cartoon characters, symbolize unity and resourcefulness, and are good at using tools to overcome any challenge they encounter.
+3.1 Framework Overview
+Figure 1 illustrates the entire workflow for the Smurfs framework.
+Initially, the Planning Agent identifies the user’s complex query and breaks it down into manageable sub-tasks.
+Executor Agents are then tasked with collecting task specific information, utilizing access to external tools.
+Answer Agent compiles the findings into a cohesive response, which is subsequently verified by the Verifier Agent to ensure accuracy and relevance.
+By dividing tasks among different agents, each agent can focus on a specific part of the task, accessing only the necessary history as context during task execution, which effectively addresses the issue of redundant context.
 The redesign of the rollback mechanism to incorporate memory and tool list rollback mechanisms addresses the instability of the rollback mechanism.
@@
 (2023), the original problem is first decomposed into sub-problems for macro-level planning.
-Subsequently, DFSDT is used to solve each sub-problem at the micro-level, with macro-level planning guiding the micro-level planning, thereby resolving the issue of premature termination.
-Figure 2: Details of the subtask-solving process of the Smurfs framework. The dotted line represents that the agent can see the memory and the full line stands for operation.
-3 Smurfs: A framework with multiple agents
-The Smurfs, the beloved cartoon characters, symbolize unity and resourcefulness, and are good at using tools to overcome any challenge they encounter.
-3.1 Framework Overview
-Figure 1 illustrates the entire workflow for the Smurfs framework.
-Initially, the Planning Agent identifies the user’s complex request and breaks it down into manageable sub-tasks.
-Executor Agents are then tasked with collecting task specific information, utilizing access to external tools.
-Answer Agent compiles the findings into a cohesive response, which is subsequently verified by the Verifier Agent to ensure accuracy and relevance.
-Each agent focus on its own task and only use the relevant part of the conversation history to reduce the Redundant Context.
-This process exemplifies the framework’s capability to efficiently handle complex queries by leveraging the specialized roles of multiple agents, thereby ensuring both the precision of task execution and the quality of the output.
-In the following sections, the system mechanism and functions of each agent will be detailed.
-More details of memory system can be seen at Appendix B.
+Subsequently, Smurfs is used to solve each sub-problem at the micro-level, with macro-level planning guiding the micro-level planning, thereby resolving the issue of premature termination.
+In the rest of this section, the mechanism of the system and the functions of each agent will be detailed. More details of memory system can be seen at Appendix C.
 3.2 Agent Components
+In this section, we introduce the two core components of the Smurfs system:
 Tools
@@
 Where q is sub-problem from the Planning Agent, r is response from the Executor Agent, M is the local memory (or global memory if max retry reaches).
-As the ‘lost-in-the-middle’ theory described in section 2.1, retaining all information may not always be beneficial, particularly in cases where the solution path is challenging to discern.
+As the ‘lost-in-the-middle’ theory described in section 1, retaining all information may not always be beneficial, particularly in cases where the solution path is challenging to discern.
 Therefore, the primary role of the Answer Agent is to succinctly summarize the generated answers and tool responses to maintain the memory efficiency.
@@
 We conducted a detailed analysis of the token costs associated with each tool planning method for the tasks, a critical evaluation aspect for multihop reasoning tasks.
-As shown in Table 1 (detailed in Appendix E), the average token costs per question and API request are evaluated for ReACT, DFSDT, and Smurfs on StableToolBench.
+As shown in Table 1 (detailed in Appendix G), the average token costs per question and API request are evaluated for ReACT, DFSDT, and Smurfs on StableToolBench.
 The analysis reveals that DFSDT generally requires about 20,000 tokens per question, encompassing both prompt and completion tokens.
@@
 The evaluation metrics is $\texttt{reward}\in[0,1]$, defined as the F1 score grading between the prediction and ground-truth answer.
-For more details about the experiment, see Appendix C.
+For more details about the experiment, see Appendix D.
 Results
@@
 This indicates that LLama-2-13b-chat may need further training for usage of tools.
-The experimental results may substantiate this viewpoint, demonstrating that the untrained methods of llama-2-13b-chat generally exhibit significantly lower accuracy compared to the trained methods.
-Nevertheless, Smurfs achieves the second highest accuracy among the untrained methods, only slightly behind reflexion, which still attests to Smurfs’ capability.
+The experiment results may substantiate this viewpoint, demonstrating that the untrained methods of llama-2-13b-chat generally exhibit significantly lower accuracy compared to the trained methods.
+To provide further insight into the model’s performance in the tool planning process, we manually categorized the types of errors made by Smurfs on the hotpotQA hard dataset in appendix E.
 I3-Inst.
@@
    w/o Planning Agent $35.5_{\pm 3.3}$ $42.6$
+   w/o Planning & Verifier Agent $58.5_{\pm 2.0}$ $57.4$
 GPT-4 Turbo with Smurfs $70.5_{\pm 1.0}$ 72.1
@@
    w/o Planning Agent $71.9_{\pm 2.8}$ $63.9$
+   w/o Planning & Verifier Agent $79.8_{\pm 2.8}$ $67.2$
 Table 5: Ablation study on StableToolBench I3-Inst subset to investigate the importance of each component within the framework.
@@
 We removed each agent individually, except for the indispensable Executor Agent, and compared the results to the complete framework.
-Table 5 shows that the Planning Agent is the most crucial component, followed by the Verifier Agent, with the Answer Agent being the least important.
-(1) Verifier Agent Removal: Without verification, the framework uses a general depth-first search, leading to increased computational demand and more tool invocations.
-(2) Answer Agent Removal: Removing this agent means the Executor Agent’s answers won’t be summarized, risking the ’lost-in-the-middle’ problem due to lengthy tool responses.
-As shown in the results, a more intelligent model, GPT-4 Turbo, can mitigate the negative impact of the Answer Agent’s removal.
-We believe this is because the more powerful model can leverage more information effectively.
-(3) Planning Agent Removal: Removing the Planning Agent affects the global path-searching strategy.
-Models with Smurfs may show reduced performance without preliminary planning, as seen in current frameworks like ReACT and DFSDT.
-The results demonstrate that the impact of removing the Planning Agent is significant, as it directly influences the multihop reasoning ability of the MAS.
+Settings
+(1) Verifier Agent Removal: Without Verifier Agent, the framework uses a generatl DFSDT in each subtask, i.e.
+including a finish tool into the tool library and deciding whether to stop at tool choice step.
+(2) Answer Agent Removal: Without Answer Agent, the framework uses full tool response instead of the summary of tool response in its memory.
+(3) Planning Agent Removal: Without Planning Agent, the framework uses Verifier Agent to decide whether the task is completed.
+(4) Planning & Verifier Agent Removal: Without Planning and Verifier Agent, the only difference between Smurfs and DFSDT will be the Answer Agent, i.e.
+including a finish tool in the tool library, deciding whether to stop at tool choice step and using summary of tool response in its memory.
+Results
+Table 5 shows the experiment result, highlighting several key insights regarding the impact of different components in the multi-agent system.
+First, the removal of any component generally reduces the win rate, highlighting each component’s significance.
+Second, performance degradation trends remain consistent: Eliminating the Answer Agent has minimal impact, whereas removing the Planning Agent causes the greatest decline.
+However, this does not imply the relative importance of these agents but rather suggests that the Verifier-Planning Agent workflow is more robust than the Verifier-Answer Agent workflow.
+Third, the performance impact varies by model capability.
+For GPT-3.5, the removal of the Answer Agent reduces both the pass and the win rates, while for GPT-4, it preserves the win rate and improves the pass rate, probably due to the superior long-text reasoning ability of GPT-4, which processes additional tool response more effectively.
+Removal of the Planning Agent significantly reduces the pass rate of GPT-3.5 but has a minimal effect on GPT-4, suggesting that the GPT-4 Verifier Agent is more robust.
+These findings may indicate that more powerful models can compensate for agent removal, sometimes even benefiting from additional context, while weaker models benefit more from complex workflow and compressed context.
+We also noticed an interesting phenomenon: when only the Planning Agent is removed, the system experiences a major decline in performance, while removing the Verifier Agent in addition to the Planning Agent improves performance again.
+This may indicate that the Verifier Agent’s ability to pause tasks or provide next-step guidance is suboptimal and requires further training.
+In conclusion, our findings offer some insight into the relationship between model capabilities and multi-agent system performance.
+The optimal number of agents and workflow may vary depending on the capacity of the model.
+We propose the hypothesis that weaker models benefit more from complex multi-agent systems and context segmentation, while stronger models perform better with comprehensive context and simpler agent systems.
+To validate this hypothesis, more extensive ablation studies are needed under a wider range of models and constraints to explore the influence of different context structures and reasoning workflows.
+We leave this investigation for future work.
 5.2 Case Study
@@
 6 Conclusion
-In this study, we present a novel MAS framework, ‘Smurfs’, tailored to enhance the planning and reasoning capabilities of LLMs in handling complex tasks that involve lengthy contexts and tools.
-We conduct experiments on the multi-step tool usage benchmark, StableToolBench and HotpotQA, and the results demonstrate the overall effectiveness and efficiency of the Smurfs framework compared to baseline methods.
-In conclusion, this research contributes to the expanding field of study focused on enhancing LLM capabilities, particularly for multi-step tool usage tasks.
-It emphasizes the importance of task decomposition, preliminary planning, and efficient verification for improving task execution performance.
-For future work, we believe incorporating more dedicated and specific roles within the system may further enhance effectiveness and efficiency, based on the ‘Smurfs principle’: synergistic collaboration among specialized agents can overcome the limitations faced by individual LLMs.
-7 Limitations
+In this paper, we propose Smurfs, an innovative MAS framework designed to enhance the tool-planning capabilities of LLMs without requiring additional training.
+Through extensive experiments on both open-ended and closed-ended tool planning benchmarks, Smurfs demonstrate its effectiveness by consistently outperforming baseline methods.
+Ablation study further provides deeper insights into the impact of each agent in our framework.
+Based on the ablation study, we propose a hypothesis that can be further tested through more comprehensive experiments in future research: weaker models benefit more from complex multi-agent systems and context segmentation, while stronger models perform better with comprehensive context and simpler agent systems.
+The findings not only advance the state-of-the-art in multi-tool planning systems but also highlight the potential of modular, training-free frameworks for LLMs in various practical applications.
+Looking forward, future research could focus on exploring Smurfs’ use in new domains, such as facilitating the synthesis of high-quality multi-tool planning data and enhancing the base model’s reasoning and tool-use abilities, further advancing the field of adaptive AI systems.
+Limitations
 Model Size Constraints:
-Due to computational constraints, our experiments did not include larger and more diverse types of LLMs.
+Due to computational constraints, our experiments did not include larger and more diverse types of LLMs. We believe this would not affect the main observations of this paper.
 Agent Component Scale-Up:
@@
 Acknowledging these limitations, future research should aim to address these gaps to provide a more comprehensive understanding of the Smurfs framework’s capabilities and potential areas for improvement.
-Appendix A Details of DFSDT
-See Figure 5.
-Figure 5: A comparison of DFSDT and conventional CoT or ReACT during model reasoning (left) Qin et al. (2024). We show part of the solution path annotation process using GPT-4 (right).
-Appendix B Details of the Smurfs
-See Figure 6 for executor working process and Figure 4 for memory and tool library of Smurfs.
+Acknowledgements
+This work was supported by the Shenzhen Science and Technology Program (JCYJ20220818103001002), Shenzhen Doctoral Startup Funding (RCBS20221008093330065), Tianyuan Fund for Mathematics of National Natural Science Foundation of China (NSFC) (12326608), Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (grant number ZDSYS20230626091302006), and Shenzhen Stability Science Program 2023.
+Appendix A Comparison between Smurfs and existing MAS
+In conclusion, Smurfs stands out compared to existing MAS in followig terms:
+Flexibility
+Smurfs utilize enhanced version of DFSDT, which intergrades global and iterative planning, while existing MAS mainly use ReAct and iterative planning.
+This make Smurfs more flexible in the planning process.
+Adaptability
+Smurfs realize superior performance on multiple different benchmarks, proving the adaptability of Smurfs.
+Existing MAS like RestGPT, Lumos are tailored for specific downstream task and need additional training to be used in other scenario.
+Learning Efficiency
+Smurfs do not need training, thus have the highest learning efficiency among existing MAS. It only need query and tool documentation from the user.
+To further illustrate these points, we conducted a detailed comparison between Smurfs and two well-known Multi-Agent Systems, highlighting their differences and the adjustments required when learning out-of-box tasks.
+CAMEL
+CAMEL Li et al.
+(2023) is a communicative agent framework.
+It uses role-play technique and inception-prompting technique to achieve autonomous cooperation between agents.
+CAMEL does not natively support tool use settings.
+CAMEL is considered to perform poorly when generalizing to new tasks (e.g., on the MATH dataset Wu et al.
+(2023)).
+Autogen
+Autogen Wu et al.
+(2023) is not designed as an agent framework for any specific task scenario.
+Instead, it provides a multi-agent conversation framework that allows users to customize agent characteristics and complete tasks through discussions among different agents.
+Autogen addresses various problems not by employing a uniform workflow, but by allowing users to design customized agents and workflows flexibly based on their tasks.
+Smurfs
+Smurfs is designed as a unified workflow specifically for complex multi-tool planning scenarios.
+For different tasks, Smurfs only needs to adjust the few-shot examples in the agent prompts and provide documentation for the tools applicable to the task, allowing Smurfs to generalize to other task scenarios.
+Smurfs was initially designed for the Stable-toolbench, which itself encompasses various types of tasks and has access to over 16,000 plugins.
+HotpotQA was subsequently introduced to evaluate the performance of Smurfs on closed-end tasks.
+When migrating from Stabletoolbench to HotpotQA, only the few-shot examples in the planning agent prompts were modified, along with the provision of plugin documentation for HotpotQA.
+The rest of the system continued using the same unified framework as in Stabletoolbench.
+Experiment results demonstrate that our untrained out-of-box unified framework achieves and even surpasses performance of agent systems specifically trained on HotpotQA, such as Autoact and Fireact, showcasing Smurfs’ flexibility in generalization.
+Compression Method Applicable Scenario Compressed Object Implementation
+Gist Tokens Mu et al. (2024) General scenarios Frequently used system prompts Training LLM to compress system prompts, reducing token usage.
+CAMPHOR Fu et al. (2024) Tool use scenarios Tool descriptions Adopts a similar approach to gist tokens, compressing each tool description into a single token, thus reducing token cost.
+Smurfs Complex multi-tool planning scenarios Input context in tool planning Operates on multi-tool planning workflows, compressing the context needed for each tool planning process.
+Table 6: Comparison of Token Compression Methods
+Appendix B Details of DFSDT
+DFSDT Qin et al.
+(2024) gives control to the model to stop and rollback the solution trajectory by using Finish tool, thus addressing limitations of ReACT.
+Finish tool has two parameters give answer: model thinks the task is finished and decide to give answer and give up and restart: model thinks current trajectory can’t lead to correct answer and decide to rollback.
+Appendix C Details of the Smurfs
+Executor Agent Details
+As illustrated in Figure 5, given a sub-problem p, Executor Agent first thinks about what to do this time, generates thought $\gamma$ according to p, local memory M, hint h from the Verifier Agent and tool list at the current step $\tau$.
+Then it will choose action $\alpha$ to complete the sub-problem using p, $\gamma$ and $\tau$.
+After that, parameters of $\alpha$ are generated using p, local memory M and tool description of the action $D[\alpha]$.
+Tool is then invoked to complete the task.
+Memory Details
+As illustrated in Figure 4, there are four kinds of memory in Smurfs.
+Local memory stores thought-answer pairs of the current solution trajectory, while global memory stores all history solution trajectory (including those that is backtracked).
+Tool list only stores available tools’ name and its usage description, while tool doc stores all detailed information about the tools including parameters details.
+Through using different kinds of memory under different circumstances, Smurfs can use DFSDT in a context efficient way.
+Restart Mechanism
+Every time Smurfs generate an intermediate output, a format checker is used to check whether the output is of the expected format.
+If not, Smurfs will retry the same step until reach retry limit or generate correct format output.
+This mechanism is used in addition to the rollback mechanism to handle the situation where the system can generate correct content but fail to follow the output format.
 Figure 5: Details of the executor agent working process
-Appendix C Experiment Settings for Hotpot QA
+Appendix D Experiment Settings for Hotpot QA
 Following settings in Qiao et al.
@@
 Table 7: Tool library for HotpotQA.
-Appendix D Prompts for multi-agent implementation
-Prompts used by each agent and their example outputs are shown in Figure  7 to  13.
-Appendix E Token Cost on StableToolBench Evaluation
+Appendix E Error Analysis on HotpotQA
+To provide further insight into the model performance in the tool planning process, we manually categorized the types of error made by Smurfs in the hotpotQA hard dataset.
+Table 8 shows that the most frequent errors committed by mistral-7b Smurfs is tool argument fail, followed by bad planning and answer miss.
+Smurfs do not make tool choice errors and premature termination errors.
+This shows that Smurfs actually alleviates the premature termination problem, making the tool choice process more robust.
+Additionally, we note that a portion of the error samples are false negatives which arise when the generated answers differ in expression from the ground truth but are equivalent in meaning.
+This highlights potential directions for future improvements in Smurfs.
+Bad Planning Answer Miss Tool Wrong Tool Argument Fail False Negative Premature Termination Total Accuracy
+Mistral-7b-Smurfs 11 7 0 28 12 0 0.42
+Table 8: Error analysis for Smurfs on HotpotQA Hard.
+Appendix F Prompts for multi-agent implementation
+Prompts used by each agent and their example outputs are shown in Figure  6 to  12.
+Appendix G Token Cost on StableToolBench Evaluation
 We analyzed the token cost for the StableToolBench experiments.